### v1

# Mean-field games with rough common noise I: the linear-quadratic case

Abstract
Motivated by mean-field games (MFG) with common noise on the one hand and pathwise stochastic control theory on the other, we formulate here a linear-quadratic (LQ) MFG with rough common noise, along with a satisfactory well-posedness theory for the linear-quadratic case. A novel Volterra-type (or mild) formulation allows to keep technical (rough-stochastic) consideration to a minimum. We derive a characterization of the optimal state and optimal control through a rough forward-backward SDE (rough FBSDE), and provide an existence and uniqueness result under the usual assumptions. Our theory is accompanied by stability estimates with respect to initial data and common noise while we also establish continuity of what we call the Itô-Lions-Lyons map for rough mean-field games. Finally, we discuss randomization of the rough common noise under appropriate conditions on the coefficients. When the latter is given by the Stratonovich lift of a Brownian motion independent of the idiosyncratic noise, we show that solutions of the rough LQ MFG coincide with those obtained by conditioning on the common noise.
Keywords. Linear-quadratic, stochastic optimal control, mean field games, rough paths, common noise, stochastic Volterra equations, Riccati equations, rough forward-backward SDEs.
MSC (2020). 91A16, 93E20, 49N10, 60L20, 60H20, 45D05, 60L50.
1 Introduction
Mean field game (MFG) theory provides a mathematical framework for analyzing strategic interactions in large populations of agents. Introduced independently by Lasry and Lions [47, 48] and Huang, Caines, and Malhamé [44, 43], MFGs describe the asymptotic behavior of Nash equilibria in $N$-player stochastic differential games as $N\to\infty$, when each individual player has negligible influence on the overall system but interacts through aggregate quantities such as the empirical distribution of states. The central idea of MFG theory is to replace the $N$-player game with a representative-agent control problem coupled to a consistency condition: each agent optimizes a cost functional that depends on the time-evolving distribution of the population and, in equilibrium, this distribution must coincide with the law of the optimally controlled state. This yields a characterization of equilibria as a fixed-point problem in the space of probability measures.
Since their introduction, MFGs have driven substantial mathematical innovation—particularly in PDE and probability theory—and have found numerous applications in engineering, economics, and operations research. In engineering, decentralized control of robotic swarms, communication networks, and automated vehicles relies on average population behavior [2, 44, 43, 10]. In finance and economics, applications range from portfolio choice problems under performance concerns [16, 38, 46] to optimal trade execution under price impact [15, 14, 36, 37], as well as optimal harvesting of exhaustible resources and models of optimal energy transition [22, 27, 53, 41, 40].
Over the past two decades, two complementary approaches to MFG theory have emerged: an analytic (PDE) approach and a probabilistic (stochastic) approach. In their original formulation [47, 48], Lasry and Lions characterized MFG equilibria through a coupled system of nonlinear partial differential equations consisting of a backward Hamilton–Jacobi–Bellman (HJB) equation for the representative agent’s value function and a forward Fokker–Planck (Kolmogorov) equation describing the evolution of the population distribution. The probabilistic approach was pioneered by Carmona and Delarue (see their two-volume monograph [19, 18]). In this framework, the representative agent’s state follows a Stochastic Differential Equation (SDE) whose coefficients depend on the law of the state process itself, and equilibria are characterized via forward–backward SDEs (FBSDEs) of McKean–Vlasov type. The probabilistic formulation is particularly powerful in the presence of common noise, where the population distribution becomes stochastic and evolves as a random measure. Recent advances in rough stochastic analysis, control, and mean field models address rough semimartingales, pathwise control and SPDEs, rough SDEs, Malliavin calculus, McKean–Vlasov BSDEs, and modern control principles [32, 30, 12, 51, 28, 21, 42, 34].
In this paper, we study a novel class of (linear–quadratic) mean field games with rough common noise. Although rough path theory has now a substantial history in applications to (stochastic) optimal control, mean-field dynamics and finance by, e.g. [20, 24, 4, 7, 8, 5, 28, 21, 9, 33, 42], our work is (to the best of our knowledge) the first to incorporate rough path techniques into the analysis of mean field games. That said, the desirability of bringing these topics together was already pointed out by Carmona–Delarue [17]:
“When analyzed within the probabilistic framework of the stochastic maximum principle, MFGs with a common noise […] best understood [with] rough paths. Indeed integrals and differentials with respect to the conditioned paths can be interpreted in the sense of rough paths while the meaning of the others can remain in the classical Itô calculus framework.”
A sufficiently flexible “hybrid” rough Itô theory arguably did not take shape until [30, 32, 12], despite related aspects in (rough) BSDE and stochastic filtering theory [25, 23, 26]; also [6, 13] for recent contributions.
In the present work, in a setting of linear but multivariate agent dynamics, we prove the existence and uniqueness of equilibria and establish a stability result of equilibrium aggregate dynamics on the rough path. This can be viewed as a first step towards a general stability result of MFG equilibria w.r.t. common noise. While our work heavily draws on recent advances in rough and stochastic analysis, as exposed in aforementioned references, its implementation remains surprisingly self-contained, due to an original mild (Volterra) formulation of the relevant dynamics.
1.1 Problem formulation
We state a general formulation of rough MFGs which we will later specialise to our linear-quadratic setting.
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, a time horizon $T$, a $q$-dimensional Brownian motion $W$, a rough path $\boldsymbol{\eta}$, and suitable functions $\ell,\Phi,b,b^{1},\Sigma$, does there exist a flow of probability measures $\{\mu_{t}\}_{t\in[0,T]}$ for which there exists
  1. (i)
    a unique minimizer $\alpha^{*}$, from some set $\mathcal{A}$ (usually accompanied with some version of convexity property to make possible the minimization), of the functional
    $\mathcal{J}(\alpha;\boldsymbol{\eta},\mu,\xi):=\mathbb{E}\Big[\int_{0}^{T}\ell(t,X^{\boldsymbol{\eta},\alpha,\mu,\xi}_{t},\mu_{t},\alpha_{t})\mathrm{d}t+\Phi(X^{\boldsymbol{\eta},\alpha,\mu,\xi}_{T},\mu_{T})\Big],$
    where
    $X^{\boldsymbol{\eta},\alpha,\mu,\xi}_{t}=\xi+\int_{0}^{t}$ $b(s,X^{\boldsymbol{\eta},\alpha,\mu,\xi}_{s},\mu_{s},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\Sigma(s,X^{\boldsymbol{\eta},\alpha,\mu,\xi},\mu_{s},\alpha_{s})\mathrm{d}W_{s}$
    $+\int_{0}^{t}b^{1}(s,X^{\boldsymbol{\eta},\alpha,\mu,\xi},\mu_{s})\mathrm{d}\boldsymbol{\eta}_{s},$
    (control problem of the representative agent)
  2. (ii)
    while also the mean field equilibrium condition holds,
    $\mu_{t}=\mathcal{L}(X^{\boldsymbol{\eta},\alpha^{*}(\mu),\mu,\xi}_{t}),\quad\alpha^{*}(\mu):=\mathrm{argmin}_{\alpha}\,\mathcal{J}(\alpha\,;\boldsymbol{\eta},\mu,\xi).$
If such $\mu$ exists it will depend on the rough driving noise $\boldsymbol{\eta}$, naturally written as $\mu^{\boldsymbol{\eta}}$, possibly multivalued. In case of uniqueness, the map from rough common noise to the equilibrium
$\boldsymbol{\eta}\mapsto\mu^{\boldsymbol{\eta}}$
is well-defined and clearly of interest, a sort of Itô-Lions-Lyons map for rough mean-field games.
We remark that even in absence of rough commen noise $\boldsymbol{\eta}$, well-posedness theory of MFG is a highly non-trival matter (e.g. [19]), with global results (in time) typically restricted to rigid structural assumptions. Fairly general results are possible for the important case of linear-quadratic mean field games (LQ MFG).
Upon randomization, say $\boldsymbol{\eta}\rightsquigarrow\boldsymbol{W}^{\perp}(\omega)$, rough path lift of some independent “common” Brownian factor $W^{\perp}$, one expects the cost functional to become
$\mathcal{J}(\alpha;\mu,\xi;\omega):=\mathbb{E}\Big[\int_{0}^{T}\ell(t,X^{\alpha,\mu,\xi}_{t},\mu_{t},\alpha_{t})\mathrm{d}t+\Phi(X^{\alpha,\mu,\xi}_{T},\mu_{T})\Big|\mathcal{F}^{\perp}_{T}\Big]\hskip 0.0pt,\\$
with “common noise” agent dynamics
$X_{t}^{\alpha,\mu,\xi}=\xi+\int_{0}^{t}b(s,X_{s}^{\alpha,\mu,\xi},\mu_{s},\alpha_{s})\hskip 1.69998pt\mathrm{d}s+\int_{0}^{t}\Sigma(s,X_{s}^{\alpha,\mu,\xi},\mu_{s},\alpha_{s})\hskip 1.69998pt\mathrm{d}W_{s}+\int_{0}^{t}b^{1}(s,X_{s}^{\alpha,\mu,\xi},\mu_{s})\hskip 1.69998pt\mathrm{d}W_{s}^{\perp},$
and equilibrium condition, for all $t\in[0,T]$,
$\mu_{t}=\mathcal{L}\left(X_{t}^{\alpha^{\ast}(\mu),\mu,\xi}\hskip 1.69998pt|\hskip 1.69998pt\mathcal{F}^{{\perp}}_{T}\right).$
Note that the control aspect of this problem corresponds precisely to the pathwise stochastic control problem of [11], first suggested by Lions-Souganidis [49, 1]. In particular, (near) optimal controls are not expected to be jointly adapted. Agent dynamics, if intrinsically formulated, would then require (partially) anticipating stochastic analysis. These difficulties evaporate when moving to a hybrid rough stochastic formulation, at the expense of working hybrid rough Itô setting, with an intrinsic formulation of the rough stochastic agent dynamics [32, 30]. For example, when we have a setup where $\Omega:=\Omega^{\prime}\times\Omega^{\prime\prime}$ that supports two independent Brownian motions $W(\omega)=W(\omega^{\prime}),{W}^{\perp}(\omega)={W}^{\perp}(\omega^{\prime\prime})$, and we define the solution of the above agent dynamics as
$\overline{X}^{\overline{\alpha}}_{t}(\omega)={X}^{\boldsymbol{\eta},{\alpha}(\boldsymbol{\eta}),\mu(\boldsymbol{\eta}),\xi}_{t}(\omega^{\prime})\big|_{\boldsymbol{\eta}={W}^{\perp}(\omega^{\prime\prime})}.$
We can expect (cf. [33, 35]) that under suitable conditions
$\mathcal{L}\left(\overline{X}^{\overline{\alpha}}_{t}\hskip 1.69998pt|\hskip 1.69998pt\mathcal{F}^{{\perp}}_{T}\right)=\mathcal{L}\left({X}^{\boldsymbol{\eta},{\alpha}(\boldsymbol{\eta}),\mu(\boldsymbol{\eta}),\xi}_{t}\right)\Big|_{\boldsymbol{\eta}={W}^{\perp}(\omega^{\prime\prime})}.$
Noting that in this situation $\mathcal{J}({\alpha};\mu,\xi;\omega)=\mathcal{J}({\alpha};\mu,\xi;\omega^{\prime\prime})$ and ${\mu}(\omega)={\mu}(\omega^{\prime\prime})$. The question arises if or when one has
$\text{essinf}_{\alpha\in\mathcal{A}}\mathcal{J}(\alpha;\mu,\xi;\omega)=\left(\inf_{\alpha\in\mathcal{A}}\mathcal{J}(\alpha;\boldsymbol{\eta},\mu,\xi)\right)\bigg|_{\boldsymbol{\eta}={W}^{\perp}(\omega^{\prime\prime})}.$
In the pure control setting of [28], [42] this was seen to hold in great generality. (We shall arrive at similar results in the case of LQ MFGs in Section 6.)
1.2 Main results
We focus on LQ MFGs where the common noise is specified by a rough path $\boldsymbol{\eta}$. More specifically, we consider a representative agent with (linear, rough) stochastic dynamics of the form
$X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}$ $=\xi+\int_{0}^{t}\big(A_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s}$ (1)
$+\int_{0}^{t}\big(A^{1}_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s},$
where $(m,m^{\prime})$ is a fixed controlled rough path and $C^{1}m\mathrm{d}\boldsymbol{\eta}$ a meaningful (Lyons-Gubinelli) rough integral (at first reading, assume $A^{1},C^{1}$ may be assumed constant in time, but we also treat controlled time-dependence, with Gubinelli derivatives $(A^{1})^{\prime},(C^{1})^{\prime}$). The agent aims to minimize the quadratic cost functional
$\mathcal{J}(\alpha)=\mathbb{E}\!\bigg[\int_{0}^{T}\frac{1}{2}\Big((X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t})^{\top}Q_{t}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}+\alpha_{t}^{\top}R_{t}\alpha_{t}\\ +(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}-S_{t}m_{t})^{\top}\bar{Q}_{t}(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}-S_{t}m_{t})\Big)\mathrm{d}t\\ +\frac{1}{2}\Big(\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}\big)^{\top}QX^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}+\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}-Sm_{T}\big)^{\top}\bar{Q}\big(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{T}-Sm_{T}\big)\Big)\bigg].$ (2)
The equilibrium condition is then, for all $t\in[0,T]$,
$m_{t}=\mathbb{E}\Big[X^{\boldsymbol{\eta},\alpha^{*},(m,m^{\prime}),\xi}_{t}\Big],\quad m^{\prime}_{t}=(A_{t}^{1}+C_{t}^{1})\mathbb{E}\Big[X^{\boldsymbol{\eta},\alpha^{*},(m,m^{\prime}),\xi}_{t}\Big].$ (3)
It is customary in LQ MFG theory, that one does not try to solve the fixed point problem with respect to the flow of probability measures directly, but rather solves the adjoint fixed point problem with respect to the corresponding mean flows. Of course the two fixed point problems are equivalent, if the initial distribution of the representative agent is Gaussian. In our setting, the mean flow formulation is more beneficial since a) it leads to a natural extension of the classical theory to a rough setting and b) it does not make use of measure-valued paths/flows and avoids associated technical considerations.
Our results can be seen as extensions of [19, Thm 3.34], also Proposition 3.35 therein, on global solutions. A novel feature or our analysis is to exploit the affine-linear rough noise structure of the agent dynamics (1) to rewrite it in a mild/Volterra form. Our first main result, stated here in a simplified form, establishes the equivalence between Volterra and rough MFGs; the precise formulation and proof follows from the combination of Theorems 4.5, 4.8 and 4.12 below.
Theorem 1.1.
The rough LQ MFG problem (1)-(3) has a unique solution if and only if the rough FBSDE (4) has a unique solution, if and only if there exists a unique solution to the Volterra FBSDE (5).
$\left\{\begin{aligned} &X^{\xi}_{t}=\xi+\int_{0}^{t}\big(A^{1}_{s}X^{\xi}_{s}+C^{1}_{s}\mathbb{E}\big[X^{\xi}_{s}\big]\big)\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad+\int_{0}^{t}\big(A_{s}X^{\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{\xi}_{s}+C_{s}\mathbb{E}\big[X^{\xi}_{s}\big]\big)\,\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{\xi}_{t}=(Q+\bar{Q})X^{\xi}_{T}-\bar{Q}S\mathbb{E}\big[X^{\xi}_{T}\big]+\int_{t}^{T}(A^{1}_{s})^{\top}Y^{\xi}_{s}\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad+\int_{t}^{T}\big((Q_{s}+\bar{Q}_{s})X^{\xi}_{s}-\bar{Q}_{s}S_{s}\mathbb{E}\big[X^{\xi}_{s}\big]+A_{s}^{\top}Y^{\xi}_{s}\big)\mathrm{d}s-\int_{t}^{T}Z^{\xi}_{s}\mathrm{d}W_{s},\end{aligned}\right$. (4)
$\left\{\begin{aligned} &{X}^{\xi}_{t}=\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(t,0)\xi+M^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})(C^{1},(C^{1})^{\prime})}\big((\mathbb{E}\big[X^{\xi}\big],(A^{1}+C^{1})\mathbb{E}\big[X^{\xi}\big]\big)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(t,s)\big(A_{s}{X}^{\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{\xi}_{s}+C_{s}\mathbb{E}\big[X^{\xi}_{s}\big]\big)\,\mathrm{d}s\\ &\quad\quad\quad\quad+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{\xi}_{t}=\big(\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(T,t)\big)^{\top}\big((Q+\bar{Q})X_{T}^{\xi}-\bar{Q}S\mathbb{E}\big[X^{\xi}_{T}\big]\big)\\ &\quad\quad+\int_{t}^{T}\big(\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(s,t)\big)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\xi}-\bar{Q}_{s}S_{s}\mathbb{E}\big[X^{\xi}_{s}\big]+A_{s}^{\top}Y^{\xi}_{s}\Big)\mathrm{d}s\\ &\quad\quad\quad\quad-\int_{t}^{T}\big(\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(s,t)\big)^{\top}Z^{\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$. (5)
Since (affine-)linear RDEs are quite well-understood, e.g. [29], its (affine-)linear solution flows offers a mild formulation of (1), in which the crucial $\mathrm{d}\boldsymbol{\eta}$ term is absorbed, essentially, in a harmless two-parameter kernel $\mathcal{K}=\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(t,s)$ multiplicative in the sense of linear operator composition, and a deterministic initial path $M^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})(C^{1},(C^{1})^{\prime})}(\cdot)_{t}$. Instead of a rough FBSDE, we are thus led to Volterra FBSDEs,
$\left\{\begin{aligned} &{X}^{m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m,\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
This mild formulation strikes a balance between (1), which requires state of the art rough stochastic analysis techniques, and a fully (Doss-Sussman type flow) transformed equation, often seen as non-intrinsic for the problem. [^1] Note that the same approach can be applied to more general drift terms. (The mild formulation may also serve as inspiration for classes of rough FBSPDEs, with additional Laplacian.)
It remains to see when such (rough or Volterra) FBSDEs are well-posed. Our second main result provides a positive answer.
Theorem 1.2.
Let $A_{t},B_{t},C_{t},R_{t},Q_{t},\bar{Q}_{t},S_{t}$ be continuous, $R_{t}$ strictly positive definite, $Q_{t},\bar{Q}_{t},Q,\bar{Q},$
$(Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}),(Q+\bar{Q}-\bar{Q}S)$ non-negative definite, $A^{1}_{t},C^{1}_{t}$, appropriately controlled by $\boldsymbol{\eta}$, $C_{t},C^{1}_{t}$ to satisfy a specific technical assumption [^2] See assumption (S4-R) or (S5-R) below., and $\xi,\Sigma_{t}$ in appropriate spaces with sufficient integrabilities. Then, the rough LQ MFG has a unique solution. Moreover the optimal control $\alpha$ has an explicit form in terms of a backward Volterra equation.
The proof involves terminal value rough Riccati equations (RRE). The latter occur as variations of RREs that were recently found in the context of robust Kalman-Bucy filtering [13] (with observations modeled as rough path). The detailed version and proof can be found in Theorem 4.12 (well-posedness of corresponding FBRDEs).
Remark 1.3.
One can relax the continuity assumptions on the non-rough coefficients in the above theorem and replace them by boundedness and Borel measurability.
Furthermore, we show stability properties of the optimal state, optimal control and the value functional with respect to the rough common noise and initial conditions. In particular we have
Theorem 1.4 (Stability).
Under appropriate regularity conditions, the equilibrium law, optimal value function and optimal controls, respectively understood as maps
$(\boldsymbol{\eta},\xi)\longmapsto\mu^{\boldsymbol{\eta},\xi}=\mathcal{L}(X^{\boldsymbol{\eta},\alpha^{*}(\boldsymbol{\eta},\xi),\xi}_{\cdot})\;,\;\;\alpha^{*}(\boldsymbol{\eta},\xi)=\mathrm{argmin}_{\alpha\in\mathcal{A}}\,\mathcal{J}(\alpha\,;\boldsymbol{\eta},(m,m^{\prime}),\xi)\;,\;\;\mathcal{J}(\alpha^{*};\boldsymbol{\eta},(m,m^{\prime}),\xi),$
are (locally Lipschitz) continuous in rough path metrics.
A detailed version of the this theorem is contained in Section 5 below as Theorem 5.3.
Finally, Section 6 is devoted to randomization, already discussed in the end of Section $1.$
1.3 Structure of the paper
After discussing frequently-used notation and some preliminaries on rough paths and rough (stochastic) integration at Section 2, we begin our approach to rough LQ MFG at Section 3. There, we develop the stochastic maximum principle in a Volterra setting by avoiding completely any mention of rough paths. This approach allows us to avoid calculations that would otherwise require strong moment conditions that in general are not possible to verify (at least in the context of Hölder rough paths). Then, at Section 4 we connect the Volterra and rough settings by delicate use of Duhamel’s principle, rigorously proved with a rough product formula. Thus, we obtain the aforementioned existence and uniqueness results for the Volterra and Rough FBSDEs, while also reducing the problem of solutions for the Volterra and rough LQ MFG problems to their deterministic counterparts. Next, we solve the deterministic forward-backward rough differential equation (FBRDE) by following the classical approach and imposing more general assumptions to account for the presence of the rough path. In Section 5, we establish stability estimates and local-Lipschitz continuity of the Itô-Lions-Lyons map with respect to rough common noise and initial conditions. Finally, Section 6 is devoted to randomization of the common noise and connection to the (conditional) stochastic problem that was discussed at the end of Section 1.1 . Auxiliary results on linear rough SDEs, accompanied by short proofs, are relegated to Appendix A.
Acknowledgments PKF, UH and ST acknowledge support from DFG CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Fields”, Project B04.
2 Preliminaries
2.1 Frequently used notation
For two matrices $A,B$ we write $A\succeq B$ (resp. $A\succ B$) to indicate that ABA-B is symmetric and non-negative (resp. positive) definite. For an $n$-tensor $A,$ we use $A^{\top}$ to denote matrix transposes when $n=2$ and (one of the) well-defined pairwise tensor transposes for $n>2.$ By slight abuse of notation, different pairwise tensor transposes will not be distinguished when they can be inferred from the context.
For every pair $n,n^{\prime}\in\mathbb{N}$ we shall denote by $\mathbb{R}^{n\times n^{\prime}}$ the space of $n\times n^{\prime}$-matrices with real entries endowed with its Euclidean norm, i.e., for $A\in\mathbb{R}^{n\times n^{\prime}}$ $|A|:=\textrm{Tr}[AA^{\top}]$, where $A^{\top}$ is the transpose of $A$. We shall frequently identify $\mathbb{R}^{n}$ with $\mathbb{R}^{n\times 1}$.
Throughout this work we fix a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in\mathbb{R}_{+}},\mathbb{P})$ which satisfies the usual conditions, $k\in(0,\infty),T\in\mathbb{R}_{+},d,q\in\mathbb{N}$ and a Euclidean space $(V,|\cdot|)$. We use ${L}(V_{1},V_{2})$ to denote the space of linear maps between Euclidean spaces $V_{1},V_{2}.$ The Borel $\sigma$-algebra on a Polish space $\mathcal{X}$ is denoted by $\mathcal{B}(\mathcal{X}).$ For our analysis we are going to need the following spaces:
$\mathbb{L}^{k}(\mathcal{F}_{t};V)$ $:=\Big\{Z:\Omega\rightarrow V:\mathcal{F}_{t}-\text{measurable and}\hskip 5.69046pt\|Z\|_{k}:=\mathbb{E}\big[|Z|^{k}\big]^{\frac{1}{k}}<\infty\Big\},$
$\mathbb{L}^{\infty}(\mathcal{F}_{t};V)$ $:=\Big\{Z:\Omega\rightarrow V:\mathcal{F}_{t}-\text{measurable and}\hskip 5.69046pt\|Z\|_{\infty}:=\mathop{\mathrm{ess\,sup}}_{\omega}|Z(\omega)|<\infty\Big\},$
$\mathbb{S}^{k}([0,T];V)$ $:=\Big\{Z:[0,T]\times\Omega\to V:\text{ prog. measurable and}\hskip 5.69046pt\|Z\|_{\mathbb{S}^{k}}:=\mathbb{E}\Big[\sup_{t\in[0,T]}|Z_{t}|^{k}\Big]^{\frac{1}{k}}<\infty\Big\},$
and
$\mathbb{H}^{k}([0,T];\mathbb{R}^{d\times q}):=\bigg\{Z$ $:[0,T]\times\Omega\to\mathbb{R}^{d\times q}:\text{ prog. measurable and}$
$\|Z\|_{\mathbb{H}^{k}}:=\mathbb{E}\bigg[\Big(\int_{0}^{T}|Z_{t}|^{2}\mathrm{d}t\Big)^{\frac{k}{2}}\bigg]^{\frac{1}{k}}<\infty\bigg\}.$
2.2 Rough path theory
Let us recall some basic results on rough paths and rough stochastic integrals. We refer to the textbook [31] for more details on rough analysis.
In what follows, $\gamma\in(\frac{1}{3},\frac{1}{2})$ is a fixed Hölder exponent corresponding to the underlying rough path. For a one-parameter function/path $\eta:[0,T]\rightarrow V,s,t\in[0,T]$ we set $\delta\eta_{s,t}:=\eta_{t}-\eta_{s}.$ The space of $V$-valued $\gamma$-Hölder continuous functions is denoted $C^{\gamma}([0,T];V)$. For any two-parameter function $A:\Delta_{[s,t]}\rightarrow V,$ where
$\Delta_{[s,t]}:=\{(u,v)\in[s,t]^{2}\penalty 10000\ |\penalty 10000\ s\leq u<v\leq t\},$
the difference operator $\delta A_{s,u,t}$ is defined by
$\delta A_{s,u,t}:=A_{s,t}-A_{s,u}-A_{u,t},\ \ \ s\leq u\leq t,$
and the $\gamma$-Hölder seminorm is denoted
$\lvert A\rvert_{\gamma,[s,t]}:=\sup_{(u,v)\in\Delta_{[s,t]}}\frac{\lvert A_{u,v}\rvert}{|v-u|^{\gamma}}.$
The set of $\gamma$-Hölder continuous two-parameter functions is denoted $C^{\gamma}_{2}([0,T];V)$. For a random two-parameter $V$-valued function $A:\Omega\times\Delta_{[s,t]}\rightarrow V$ we set
$\lVert A\rVert_{\gamma,k,[s,t]}:=\sup_{(u,v)\in\Delta_{[s,t]}}\frac{\lVert A_{u,v}\rVert_{k}}{\lvert v-u\rvert^{\gamma}}.$
The space of all functions for which the above seminorm is finite is denoted $C^{\gamma}_{2}([0,T];\mathbb{L}^{k})$ and $C^{\gamma}_{2}\mathbb{L}^{k}$ denotes the subspace of all $\mathcal{F}\otimes\mathcal{B}(\Delta_{[0,T]})/\mathcal{B}(V)$-measurable elements of $C^{\gamma}_{2}([0,T];\mathbb{L}^{k})$.
In the following, we usually omit the time interval in our notations and norms if $[s,t]=[0,T]$ and write $\lesssim_{\theta}$’ to indicate that an inequality holds up to a generic constant that depends on a parameter $\theta$.
Remark 2.1.
For every $\gamma\in(0,1]$ and Banach space $V$, the space $C^{\gamma}([0,T];V)$ equipped with the norm $\|\cdot\|_{\gamma}:=\|\cdot\|_{\infty;[0,T]}+|\cdot|_{\gamma,[0,T]}$ is a Banach space.
Definition 2.2.
  • $(1)$
    A pair of (deterministic) continuous functions $\boldsymbol{\eta}:=(\eta,\eta^{(2)}):[0,T]\times\Delta_{[0,T]}\rightarrow\mathbb{R}^{p}\times\mathbb{R}^{p\times p}$ is called a $\gamma$-Hölder continuous rough path, if the following two conditions are satisfied.
[…]
    The space of $\gamma$-Hölder continuous rough paths is denoted $\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p})$ and is equipped with the following seminorm and rough path metric: for any $\boldsymbol{\eta}=(\eta,\eta^{(2)}),\bar{\boldsymbol{\eta}}=(\bar{\eta},{\bar{\eta}}^{(2)})\in\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p}),$
[…]
  • $(2)$
    $\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$ is the space of $\gamma$-Hölder geometric rough paths, defined as the completion of canonically lifted smooth paths (i.e. if $\eta$ is smooth, $\boldsymbol{\eta}:=(\eta,\int\eta d\eta)$ is the canonical lift of $\eta$) under the rough path metric $\rho_{\gamma}$. Note that $\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$ is Polish and $\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})\subset\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p})$ (see e.g. [31] for further details).
Remark 2.3.
We shall always endow $\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$with the rough path metric $\rho_{\gamma}$. Moreover, we shall frequently write $\rho_{\gamma}(\boldsymbol{\eta})$ instead of $\rho_{\gamma}(\boldsymbol{\eta},0)$.
To introduce stochastic controlled rough paths and to define rough integrals, we introduce for any integrable random two-parameter process $A$ the two-parameter conditional expectations process $\mathbb{E}_{\bullet}A$ by
$(s,t;\omega)\mapsto\mathbb{E}[A_{s,t}|\mathcal{F}_{s}](\omega)=:\mathbb{E}_{s}[A_{s,t}](\omega).$
Definition 2.4 ([30, Definition 3.1]).
Let $\boldsymbol{\eta}\in\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p})$ be a rough path.
  • $(1)$.
    A pair $(Z,Z^{\prime})$ is called a $\eta$-stochastic controlled rough path of $k$-integrability and $(\beta,\beta^{\prime})$-Hölder regularity with $\beta,\beta^{\prime}\in(0,\gamma],$ if the following hold:
    1. $(\text{i})$
      The process $Z=(Z_{t})_{t\in[0,T]}$ is $V$-valued, progressively measurable and
      $\|Z_{0}\|_{k}<\infty\quad\text{and}\quad\|\delta Z\|_{\beta,k}:=\sup_{0\leq s<t\leq T}\frac{\|\delta Z_{s,t}\|_{k}}{|t-s|^{\beta}}<\infty.$
    2. $(\text{ii})$
      The process $Z^{\prime}=(Z^{\prime}_{t})_{t\in[0,T]}$ takes values in ${L}(\mathbb{R}^{p},V)$ (throughout this work the codomain is finite-dimensional and thus we will not distinguish operator norms in the notation), it is progressively measurable and
      $\|Z^{\prime}_{0}\|_{k}<\infty\quad\text{and}\quad\|\delta Z^{\prime}\|_{\beta^{\prime},k}:=\sup_{0\leq s<t\leq T}\frac{\|\delta Z^{\prime}_{s,t}\|_{k}}{|t-s|^{\beta^{\prime}}}<\infty.$
    3. $(\text{iii})$
      The two-parameter process
      $R^{Z}_{s,t}:=\delta Z_{s,t}-Z^{\prime}_{s}\delta\eta_{s,t},\quad(s,t)\in\Delta_{[0,T]}$
      satisfies
      $\|\mathbb{E}_{\bullet}R^{Z}\|_{\beta+\beta^{\prime},k}:=\sup_{0\leq s<t\leq T}\frac{\|\mathbb{E}_{s}R^{Z}_{s,t}\|_{k}}{|t-s|^{\beta+\beta^{\prime}}}<+\infty.$
    In this case we write $(Z,Z^{\prime})\in\textbf{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{k}([0,T];V)$ and $(Z,Z^{\prime})\in\mathbf{D}_{\eta}^{2\beta}\mathbb{L}^{k}([0,T];V)$ if $\beta=\beta^{\prime}$. We equip the space $\mathbf{D}_{\eta}^{\beta,\beta^{\prime}}L^{k}$ with the norm
    $\llbracket(Z,Z^{\prime})\rrbracket_{\eta,\beta,\beta^{\prime},k}:=\left\lVert Z_{0}\right\rVert_{k}+\lVert\delta Z\rVert_{\beta,k}+\left\lVert Z^{\prime}_{0}\right\rVert_{k}+\lVert\delta Z^{\prime}\rVert_{\beta^{\prime},k}+\lVert\mathbb{E}_{\bullet}R^{Z}\rVert_{\beta+\beta^{\prime},k}.$
    Moreover, for any $\boldsymbol{\eta},\bar{\boldsymbol{\eta}}\in\mathscr{C}^{\gamma},$ and pairs $(Z,Z^{\prime})\in\textbf{D}^{\beta,\beta^{\prime}}_{\eta}\mathbb{L}^{k},$ $(\bar{Z},\bar{Z}^{\prime})\in\textbf{D}^{\beta,\beta^{\prime}}_{\bar{\eta}}\mathbb{L}^{k}$ we define the distance
    $\llbracket Z,Z^{\prime};\bar{Z},\bar{Z}^{\prime}\rrbracket_{\eta,\bar{\eta},\beta,\beta^{\prime},k}:=\|\delta(Z-\bar{Z})\|_{\beta,k}+\|\delta(Z^{\prime}-\bar{Z}^{\prime})\|_{\beta^{\prime},k}+\|Z_{0}-\bar{Z}_{0}\|_{k}$
    $\hskip 142.26378pt+\|Z^{\prime}_{0}-\bar{Z}^{\prime}_{0}\|_{k}+\|E_{\bullet}R^{Z}-E_{\bullet}\bar{R}^{\bar{Z}}\|_{\beta+\beta^{\prime},k},$
    where
    $\bar{R}^{\bar{Z}}_{s,t}:=\delta\bar{Z}_{s,t}-\bar{Z}^{\prime}_{s}\delta\bar{\eta}_{s,t},\quad s,t\in\Delta_{[0,T]}.$
  • $(2)$
    For any pair $(Z,Z^{\prime})\in\mathbf{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{k},$ if $R^{Z}\in C^{\beta+\beta^{\prime}}_{2}\mathbb{L}^{k}$, we write $(Z,Z^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\mathbb{L}^{k}$, and define for any $(\bar{Z},\bar{Z}^{\prime})\in\mathscr{D}_{\bar{\eta}}^{\beta,\beta^{\prime}}\mathbb{L}^{k}$, the distance
    $\|Z,Z^{\prime};\bar{Z},\bar{Z}^{\prime}\|_{\eta,\bar{\eta},\beta,\beta^{\prime},k}$ $\penalty 10000\ :=\|\delta(Z-\bar{Z})\|_{\beta,k}+\|\delta(Z^{\prime}-\bar{Z}^{\prime})\|_{\beta^{\prime},k}$
    $\quad\quad+\|Z_{0}-\bar{Z}_{0}\|_{k}+\|Z^{\prime}_{0}-\bar{Z}^{\prime}_{0}\|_{k}+\|R^{Z}-\bar{R}^{\bar{Z}}\|_{\beta+\beta^{\prime},k}.$
Remark 2.5.
We emphasize that in the case where $(Z,Z^{\prime})\in\mathbf{D}_{{\eta}}^{\beta,\beta^{\prime}}\mathbb{L}^{k}$ is deterministic then immediately we get $(Z,Z^{\prime})\in\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{\infty}.$ The space of deterministic controlled rough paths is denoted by $\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}$, or $\mathscr{D}_{\eta}^{\beta,\beta^{\prime}}([0,T];V)$ when we need to emphasize the codomain $V.$ Furthermore, in the deterministic setting we abbreviate the notation and we simply write $\|\cdot,\cdot;\cdot,\cdot\|_{\eta,\bar{\eta},\beta,\beta^{\prime}}$.
Since the coefficients $A^{1},C^{1}$ in the state dynamics (1) are controlled rough paths, we frequently need to deal with products of (stochastic) controlled rough paths. The following lemma is implicitly used in several places and establishes useful estimates for such products. The proof follows immediately from the definitions.
Lemma 2.6.
Let $(Z,Z^{\prime})\in\textbf{D}^{\beta,\beta^{\prime}}_{\eta}\mathbb{L}^{\infty}$ and $(z,z^{\prime})\in\textbf{D}^{\beta,\beta^{\prime}}_{\eta}\mathbb{L}^{k}$ for some $k\in[2,\infty]$. Then
$(Zz,(Zz)^{\prime}):=(Zz,Z^{\prime}z+Zz^{\prime})\in\textbf{D}^{\beta,\beta^{\prime}}_{\eta}\mathbb{L}^{k}.$
Moreover, if
$\mathop{\mathrm{ess\,sup}}_{t,\omega}|Z_{t}|\vee\|(Z,Z^{\prime})\|_{\boldsymbol{\eta},\beta,\beta^{\prime},\infty}\leq K\quad\mbox{and}\quad\|z_{0}\|_{k}\vee\|z^{\prime}_{0}\|_{k}\leq K_{0}$
then, for some implied constants that depend increasingly on $T$,
$\|Zz\|_{\beta,k}$ $\lesssim K\|z\|_{\beta,k}+KK_{0},$
$\|(Zz)^{\prime}\|_{\beta^{\prime},k}$ $\lesssim K(\|z\|_{\beta,k}+\|z^{\prime}\|_{\beta^{\prime},k}+K_{0}),$
$\|R^{Zz}\|_{\beta,k}$ $\lesssim K(\|z\|_{\beta,k}+\|z^{\prime}\|_{\beta^{\prime},k}+K_{0})(1\vee\|\eta\|_{\beta}),$
$\|\mathbb{E}_{\bullet}R^{Zz}\|_{\beta+\beta^{\prime},k}$ $\lesssim K(\|z\|_{\beta,k}+\|\mathbb{E}_{\bullet}R^{z}\|_{\beta+\beta^{\prime},k}+K_{0}),$
where $a\vee b:=\max{(a,b)}$.
Having introduced rough paths, we proceed to introduce rough stochastic integrals.
Proposition 2.7 ([30, Theorem 3.4]).
Let $p\in\mathbb{N},$ $\boldsymbol{\eta}\in\mathscr{C}^{\gamma}([0,T];\mathbb{R}^{p})$ a rough path, with $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. For any $k\in[2,\infty)$ and any $\boldsymbol{\eta}$-stochastic controlled rough path $(Z,Z^{\prime})\in\textbf{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{k}$, the following rough stochastic integral is well-defined:
$\int_{0}^{t}Z_{s}\mathrm{d}\boldsymbol{\eta}_{s}:=\int_{0}^{t}(Z_{s},Z^{\prime}_{s})\mathrm{d}\boldsymbol{\eta}_{s}:=\lim_{|\Pi|\rightarrow 0}\sum_{[u,v]\in\Pi}(Z_{u}\delta\eta_{u,v}+Z^{\prime}_{u}\eta^{(2)}_{u,v}),\quad t\in[0,T].$
Here, $\Pi$ is any partition of $[0,t]$ with mesh $|\Pi|$ and the limit is taken in the sense of convergence in probability. Moreover, for any $(s,t)\in\Delta,$ we have
[…]
In particular,
$(X,X^{\prime}):=\Big(\int_{0}^{.}Z_{s}\mathrm{d}\boldsymbol{\eta}_{s},Z\Big)\in\textbf{D}^{\gamma,\beta}_{\eta}\mathbb{L}^{k}.$
Due to our hybrid rough stochastic setting, we interpret the state equation (1) as a Rough Stochastic Differential Equation (RSDE) per [30]. We now define what we mean by solutions of (1).
Definition 2.8.
Let $p\in\mathbb{N},k\in[2,\infty),$ $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$, and $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. Additionally, let $A,B,C:[0,T]\rightarrow\mathbb{R}^{d\times d}$ bounded and Borel measurable on $[0,T]$, $\big({A}^{1},(A^{1})^{\prime}\big),\big({C}^{1},(C^{1})^{\prime}\big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big),(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ deterministic controlled rough paths with respect to $\eta$, $\xi\in\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, and $\alpha(\omega,t),\Sigma(\omega,t)$ progressively measurable processes such that $\sup_{t\in[0,T]}\mathbb{E}[|\Sigma_{t}|^{k}]+\sup_{t\in[0,T]}\mathbb{E}[|\alpha_{t}|^{k}]<\infty$. We shall call a continuous adapted process $X$ an $k$-integrable solution to the RSDE (1) in the space $\textbf{D}_{\eta}^{\beta,\beta^{\prime}}\mathbb{L}^{k}([0,T];\mathbb{R}^{d})$, if the following holds:
  1. (i)
    The integral $\int_{0}^{t}\big(A_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\mathrm{d}s$ is finite a.s..
  2. (ii)
    We have
    $\big(A^{1}{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}+C^{1}m,(A^{1})^{\prime}{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}+A^{1}(A^{1}{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}+C^{1}m)+(C^{1})^{\prime}m+C^{1}m^{\prime}\big)$
    $\in\textbf{D}^{\beta,\beta^{\prime}}_{\eta}\mathbb{L}^{k}\big([0,T];\mathbb{R}^{d}\big).$
  3. (iii)
    The following integral equation holds $\mathbb{P}$-a.s. for all $t\in[0,T]$:
    $X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}$ $=\xi+\int_{0}^{t}\big(A_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s}$
    $+\int_{0}^{t}\big(A^{1}_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}.$
The next result is from [12, Thm 3.11] in case $\beta^{\prime}=\beta$, and [42, Thm 3.5] otherwise. It establishes well-posedness of (1).
Theorem 2.9.
Let $p\in\mathbb{N},$ $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$,
and $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. Additionally, let $A,B,C:[0,T]\rightarrow\mathbb{R}^{d\times d}$ bounded and Borel measurable on $[0,T]$, $\big({A}^{1},(A^{1})^{\prime}\big),\big({C}^{1},(C^{1})^{\prime}\big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big),(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ deterministic controlled rough paths with respect to $\eta$}, $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, and $\alpha(\omega,t),\Sigma(\omega,t)$ progressively measurable processes such that $\sup_{t\in[0,T]}\mathbb{E}[|\Sigma_{t}|^{k}]+\sup_{t\in[0,T]}\mathbb{E}[|\alpha_{t}|^{k}]<\infty$ for every $k\in[2,\infty)$. Then, there exists a unique solution $X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ of
$X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}$ $=\xi+\int_{0}^{t}\big(A_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s}$
$+\int_{0}^{t}\big(A^{1}_{s}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}$
with $(X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi},A^{1}X^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}+C^{1}m)\in\bigcap_{k=2}^{\infty}\textbf{D}^{\gamma,\beta}_{\eta}\mathbb{L}^{k}\big([0,T];\mathbb{R}^{d}\big)$.
3 A linear-quadratric Volterra stochastic control problem
This section is devoted to the study of an LQ stochastic control problem with stochastic Volterra state dynamics. The Volterra kernel $\mathcal{K}$ is regular and chosen to satisfy a certain flow property which appears naturally in the rough LQ MFG setting of Section 4; see also Remark 4.4 below. Throughout this section, we treat the flow $m$ as a parameter i.e. a fixed continuous path.
3.1 Assumptions, state dynamics and cost functional
From this point on (and throughout the rest of the paper), we assume that $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in\mathbb{R}_{+}},\mathbb{P})$ is the usual augmentation of a filtration generated by a $q$-dimensional Brownian motion $W$ and a $\sigma$-algebra $\mathcal{F}_{0}^{0}$ independent of WW444this assumption is made to avoid technicalities with martingale representation arguments that are used in the derivation of the Volterra FBSDEs. From now on and for the rest of this section, we will work under the following assumptions:
  1. (S1)
    We are given $\mathcal{K}:[0,T]\times[0,T]\rightarrow\mathbb{R}^{d\times d}$ with the following properties:
    1. (i)
      For every $t\in[0,T]$ we have $\mathcal{K}(t,t)=\text{Id}_{d\times d}$.
    2. (ii)
      For every $s,u,t\in[0,T]$ we have $\mathcal{K}(t,s)=\mathcal{K}(t,u)\mathcal{K}(u,s)$.
    3. (iii)
      The process $\mathcal{K}(t,0):[0,T]\rightarrow\mathbb{R}^{d\times d}$ is continuous.
  2. (S2)
    Next, let $A,B,C:[0,T]\rightarrow\mathbb{R}^{d\times d}$ continuous on $[0,T]$, and $M:C([0,T];\mathbb{R}^{d})\rightarrow C([0,T];\mathbb{R}^{d})$. Furthermore, let us also fix a $\Sigma\in\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$.
  3. (S3)
    Finally, let $Q,\bar{Q}\in\mathbb{R}^{d\times d}$ with $Q,\bar{Q}\succeq 0$, and $S_{t},Q_{t},\bar{Q}_{t},:[0,T]\rightarrow\mathbb{R}^{d\times d},R_{t}:[0,T]\rightarrow\mathbb{R}^{\kappa\times\kappa},$ continuous on $[0,T]$, with $R_{t}\succ\lambda\text{Id}_{d\times d}$, for some $\lambda>0$, and $Q_{t},\bar{Q}_{t}\succeq 0$.
Remark 3.1.
From (S1)(i),(ii) it follows that $\mathcal{K}(t,0)^{-1}=\mathcal{K}(0,t)$, for every $t\in[0,T]$.
For $\kappa\in\mathbb{N},$ $\widehat{\mathcal{A}}:=\mathbb{H}^{2}([0,T];\mathbb{R}^{\kappa})$ denotes the admissible control space, equipped with the norm $\|\cdot\|_{\mathbb{H}^{2}}.$ For every $\alpha\in\widehat{\mathcal{A}},m\in C([0,T];\mathbb{R}^{d})$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d})$, a representative agent controls a process $X^{\alpha,m,\xi}:\Omega\times[0,T]\rightarrow\mathbb{R}^{d}$ evolving as
${X}^{\alpha,m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{\alpha,m,\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s}.$ (6)
Remark 3.2.
We define $\widehat{X}^{\alpha,m,\xi}_{t}:=\mathcal{K}(0,t)\Big({X}^{\alpha,m,\xi}_{t}-M(m)_{t}\Big),t\in[0,T]$. Then equation (6) becomes equivalent to
$\widehat{X}^{\alpha,m,\xi}_{t}=\xi+\int_{0}^{t}\mathcal{K}(0,s)\big(A_{s}\mathcal{K}(s,0)\widehat{X}^{\alpha,m,\xi}_{s}+B_{s}\alpha_{s}+A_{s}M(m)_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(0,s)\Sigma_{s}\mathrm{d}W_{s}.$
The above is a classical forward affine SDE so we know that it has a unique (global) solution $\widehat{X}^{\alpha,m,\xi}\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$, see for example [19, Theorem 4.21] for an even stronger result. From this we deduce that (6) has a unique solution ${X}^{\alpha,m,\xi}=\mathcal{K}(t,0)(\widehat{X}^{\alpha,m,\xi}+M(m))\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$, for every $\alpha\in\widehat{\mathcal{A}},m\in C([0,T];\mathbb{R}^{d})$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d})$.
For a fixed path $m\in C([0,T];\mathbb{R}^{d})$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d})$, the agent seeks to minimize the quadratic cost functional
$\mathcal{J}^{m,\xi}(\alpha)=\mathbb{E}\!\left[\int_{0}^{T}\ell^{m}_{t}(X_{t}^{\alpha,m,\xi},\alpha_{t})\mathrm{d}t+\Phi^{m}(X_{T}^{\alpha,m,\xi})\right],$ (7)
over all admissible controls $\alpha\in\mathcal{A},$ where
$\ell^{m}_{t}(x,\alpha)$ $:=\frac{1}{2}\Big(x^{\top}Q_{t}x+\alpha^{\top}R_{t}\alpha+(x-S_{t}m_{t})^{\top}\bar{Q}_{t}(x-S_{t}m_{t})\Big),$
$\Phi^{m}(x)$ $:=\frac{1}{2}\Big(x^{\top}Qx+(x-Sm_{T})^{\top}\bar{Q}(x-Sm_{T})\Big).$
Notice that this cost functional coincides with the original functional (2) with the only difference that the latter also depends on an underlying rough path.
Remark 3.3.
Linear-quadratic stochastic optimal control problems for more general classes of stochastic Volterra processes (but without a mean field component) have been considered e.g. in [52]. The approach followed therein involves so called Markovian lifts of Volterra processes to infinite-dimensional spaces; such techniques arise in the setting of singular (matrix-valued) kernels (see e.g. [39] and references therein) and allow access to classical Markovian tools. In contrast, the kernels $\mathcal{K}$ appearing in our Volterra formulation of the rough MFG are quite well-behaved. Indeed, the flow property (S1) guarantees that $\mathcal{K}$ is finite on the diagonal and $\mathcal{K}(0,t)$ is an invertible matrix for each $t.$ This, in turn, allows us to solve the Volterra LQ problem by transforming the state dynamics into a classical (semimartingale-type) form and without resorting to infinite-dimensional arguments.
3.2 The stochastic maximum principle
The main result of this subsection, a stochastic maximum principle, is given in Theorem 3.7 and is preceded by several preparatory arguments. Well-posedness of the associated Volterra FBSDEs is postponed to Section 3.3.
In order to solve the stochastic optimization problem from Section 3.1, we follow the classical approach of finding a critical point $\alpha^{*}$ of $\widehat{\mathcal{A}}$ by studying the null-set of the Gâteaux derivative.
To this end, let $X,H\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$ and $\alpha,h\in\widehat{\mathcal{A}}$. From (S3) we have
$(X_{t}+H_{t})^{\top}Q_{t}(X_{t}+H_{t})-X_{t}^{\top}Q_{t}X_{t}$ $=X_{t}^{\top}Q_{t}H_{t}+H_{t}^{\top}Q_{t}X_{t}+H_{t}^{\top}Q_{t}H_{t}$
$=2H_{t}^{\top}Q_{t}X_{t}+H_{t}^{\top}Q_{t}H_{t}$
and similarly
$(\alpha_{t}+h_{t})R_{t}(\alpha_{t}+h_{t})-\alpha_{t}R_{t}\alpha_{t}$ $=2h_{t}^{\top}R_{t}\alpha_{t}+h_{t}^{\top}R_{t}h_{t},$
$(X_{t}+H_{t}-S_{t}m_{t})\bar{Q}_{t}(X_{t}+H_{t}-S_{t}m_{t})$ $-(X_{t}-S_{t}m_{t})^{\top}\bar{Q}_{t}(X_{t}-S_{t}m_{t})$
$=2H^{\top}\bar{Q}_{t}(X_{t}-S_{t}m_{t})+H_{t}^{\top}\bar{Q}_{t}H_{t}.$
Hence, with the same computations for $\Phi^{m}$, we get
$\ell^{m}_{t}(X+H,\alpha+h)-\ell^{m}_{t}(X,\alpha)$
$=H_{t}^{\top}\big[(Q_{t}+\bar{Q}_{t})X_{t}-\bar{Q}_{t}(S_{t}m_{t})\big]+H_{t}^{\top}\Big(\frac{Q_{t}+\bar{Q}_{t}}{2}\Big)H_{t}+h_{t}^{\top}R_{t}\alpha_{t}+h_{t}^{\top}\frac{R_{t}}{2}h_{t}$
                 and
$\Phi^{m}(X+H)-\Phi^{m}(X)=H_{T}^{\top}\big[(Q+\bar{Q})X_{T}-\bar{Q}(Sm_{T})\big]+H_{T}^{\top}\Big(\frac{Q+\bar{Q}}{2}\Big)H_{T}.$
To finish our preparations, we define
$R^{m}_{1}(X,H,\alpha,h):=$
$\mathbb{E}\bigg[\int_{0}^{T}\Big(H_{t}^{\top}\big[(Q_{t}+\bar{Q}_{t})X_{t}-\bar{Q}_{t}(S_{t}m_{t})\big]\Big)\mathrm{d}t+H_{T}^{\top}\big[(Q+\bar{Q})X_{T}-\bar{Q}(Sm_{T})\big]+\int_{0}^{T}(h_{t}^{\top}R_{t}\alpha_{t})\mathrm{d}t\bigg]$ (8)
and
$R^{m}_{2}(H,h):=\mathbb{E}\bigg[\int_{0}^{T}\Big(H_{t}^{\top}\Big(\frac{Q_{t}+\bar{Q}_{t}}{2}\Big)H_{t}+h_{t}^{\top}\frac{R_{t}}{2}h_{t}\Big)\mathrm{d}t+H_{T}^{\top}\Big(\frac{Q+\bar{Q}}{2}\Big)H_{T}\bigg],$ (9)
so
$\mathbb{E}\!\left[\int_{0}^{T}\ell^{m}_{t}(X_{t}+H_{t},\alpha_{t}+h_{t})\mathrm{d}t+\Phi^{m}(X_{T}+H_{T})\right]-\mathbb{E}\!\left[\int_{0}^{T}\ell^{m}_{t}(X_{t},\alpha_{t})\mathrm{d}t+\Phi^{m}(X_{T})\right]$
$=R^{m}_{1}(X,H,\alpha,h)+R^{m}_{2}(H,h).$ (10)
Now, we fix $\alpha,h\in\widehat{\mathcal{A}}$ an admissible control and admissible variation respectively. For $\varepsilon>0$, define the perturbed control $\alpha^{\varepsilon}:=\alpha+\varepsilon h$ and the corresponding state $X^{\alpha^{\varepsilon},m,\xi}$ solving (6). Since the dependence of (6) on $(x,\alpha)$ is affine, we immediately get
$\frac{X^{\alpha^{\varepsilon},m,\xi}_{t}-X^{\alpha,m,\xi}_{t}}{\varepsilon}=\int_{0}^{t}\mathcal{K}(t,s)\Big(A_{s}\frac{X^{\alpha^{\varepsilon},m,\xi}_{s}-X^{\alpha,m,\xi}_{s}}{\varepsilon}+B_{s}h_{s}\Big)\,\mathrm{d}s,$
the right-hand side above is independent of $\varepsilon$, hence the Gâteaux derivative
${(\delta X)}_{t}^{\alpha,h}:=\lim_{\varepsilon\rightarrow 0}\frac{X^{\alpha^{\varepsilon},m,\xi}_{t}-X^{\alpha,m,\xi}_{t}}{\varepsilon},$
exists and satisfies the affine variational equation
${(\delta X)}_{t}^{\alpha,h}=\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{(\delta X)}_{s}^{\alpha,h}+B_{s}h_{s}\big)\,\mathrm{d}s.$ (11)
Therefore, the map $\alpha\mapsto X^{\alpha,m,\xi}$ is Gâteaux differentiable. Moreover, notice that, again due to the affine dynamics of (6), ${(\delta X)}_{t}^{\alpha,h}$ does not depend on $\alpha$ either. Hence the notation ${(\delta X)}_{t}^{h}$ is justified and for every $\varepsilon>0,$ $h\in\widehat{\mathcal{A}}$, we have
$X^{\alpha+\varepsilon h,m,\xi}_{t}=X^{\alpha,m,\xi}_{t}+\varepsilon{(\delta X)}_{t}^{h}.$
Furthermore, $\mathcal{K}(t,s),A_{s},B_{s}$ are uniformly bounded over $s,t\in[0,T]$, hence from Tonelli’s theorem and Grönwall’s and Jensen’s inequalities, we also get the energy estimate
$|{(\delta X)}_{t}^{h}|^{2}\leq 2T^{2}\int_{0}^{t}|\mathcal{K}(t,s)|^{2}\big(|A_{s}|^{2}|{(\delta X)}_{s}^{h}|^{2}+|B_{s}|^{2}|h_{s}|^{2}\big)\,\mathrm{d}s$
$\Longrightarrow\sup_{u\in[0,t]}\mathbb{E}\Big[\big|{(\delta X)}_{u}^{h}\big|^{2}\Big]\leq C_{T}\|h\|_{\mathcal{A}}^{2}+C_{T}\int_{0}^{t}\mathbb{E}\Big[\big|{(\delta X)}_{s}^{h}\big|^{2}\Big]\mathrm{d}s$
$\Longrightarrow\sup_{t\in[0,T]}\mathbb{E}\Big[\big|{(\delta X)}_{t}^{h}\big|^{2}\Big]\leq C^{\prime}_{T}\|h\|_{\mathbb{H}^{2}}^{2},$ (12)
for some constants $C_{T},C^{\prime}_{T}$ depending on $\mathcal{K},A,B,T$.
We can now move forward to compute the Gâteaux derivative of $\mathcal{J}^{m,\xi}$ with respect to $\widehat{\mathcal{A}}$. So first, we define $H^{\varepsilon}_{t}:=\varepsilon{(\delta X)}_{t}^{h}\$and by (8) and (9) we get
$\mathcal{J}^{m,\xi}(\alpha+\varepsilon h)-\mathcal{J}^{m,\xi}(\alpha)$
$=R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,\varepsilon h)+R^{m}_{2}((\delta X)^{h},\varepsilon h)$
$=\varepsilon R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,h)+\varepsilon^{2}R^{m}_{2}((\delta X)^{h},h).$
In view of (12)
$\big|R^{m}_{2}((\delta X)^{h},h)\big|\leq C^{\prime\prime}_{T}\|h\|_{\mathbb{H}^{2}}^{2},$
for some constant $C^{\prime\prime}$ depending on $\mathcal{K},A,B,T,Q_{t},\bar{Q}_{t},R_{t},Q,\bar{Q}.$ Hence
$(\delta\mathcal{J})^{m,\xi,h}(\alpha)$ $:=\lim_{\varepsilon\rightarrow 0}\frac{\mathcal{J}^{m,\xi}(\alpha+\varepsilon h)-\mathcal{J}^{m,\xi}(\alpha)}{\varepsilon}$
$\hskip 3.41418pt=R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,h)$
$\hskip 3.41418pt=\mathbb{E}\bigg[\int_{0}^{T}\big((\delta X)^{h}_{t}\big)^{\top}\big[(Q_{t}+\bar{Q}_{t})X^{\alpha,m,\xi}_{t}-\bar{Q}_{t}(S_{t}m_{t})\big]\mathrm{d}t$
$\qquad+\big((\delta X)^{h}_{T}\big)^{\top}\big[(Q+\bar{Q})X^{\alpha,m,\xi}_{T}-\bar{Q}(Sm_{T})\big]+\int_{0}^{T}(\alpha_{t}^{\top}R_{t}h_{t})\mathrm{d}t\bigg],$ (13)
which proves Gâteaux differentiability of $\mathcal{J}^{m,\xi}(\cdot)$ with derivative $(\delta\mathcal{J})^{m,\xi,h}(\cdot)$.
We now eliminate the ${(\delta X)}^{h}$ terms above using an integration by parts argument. Let $Y^{\alpha,m,\xi}$ be an adapted process satisfying the stochastic backward (adjoint) Volterra differential equation
$Y^{\alpha,m,\xi}_{t}$ $=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)$
$\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s$
$\quad-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.$
Remark 3.4.
By defining $\widehat{Y}^{\alpha,m,\xi}_{t}:=\mathcal{K}(t,0)^{\top}{Y}^{\alpha,m,\xi}_{t}$ and $\widehat{Z}^{\alpha,m,\xi}_{t}:=\mathcal{K}(t,0)^{\top}Z^{\alpha,m,\xi}_{t}$, equation (3.2) becomes equivalent to
$\widehat{Y}^{\alpha,m,\xi}_{t}=\mathcal{K}(T,0)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)$
$\quad+\int_{t}^{T}\mathcal{K}(s,0)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathcal{K}(0,s)^{\top}\widehat{Y}^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\widehat{Z}^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.$
The latter is a classical affine backward SDE with a unique solution $(\widehat{Y}^{\alpha,m,\xi},\widehat{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$, see for example [19, Theorem 4.23], hence, for every $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0}),\alpha\in\widehat{\mathcal{A}}$ and $m\in C([0,T];\mathbb{R}^{d})$, (3.2) also has a unique solution $({Y}^{\alpha,m,\xi},{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$.
Consider the scalar process $(Y^{\alpha,m,\xi}_{t})^{\top}{(\delta X)}_{t}^{h}$. The integration-by-parts formula for semimartingales gives
$(Y^{\alpha,m,\xi}_{T})^{\top}$ ${(\delta X)}_{T}^{h}=\big(\mathcal{K}(T,0)^{\top}Y^{\alpha,m,\xi}_{T}\big)^{\top}\big(\mathcal{K}(0,T){(\delta X)}_{T}^{h}\big)$
$=\int_{0}^{T}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)^{\top}\mathrm{d}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)+\int_{0}^{T}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)^{\top}\mathrm{d}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big).$
By (11) and (3.2) we have
$\int_{0}^{T}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)^{\top}\mathrm{d}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)=\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}A_{t}{(\delta X)}_{t}^{h}\Big)\mathrm{d}t+\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}B_{t}h_{t}\Big)\mathrm{d}t,$
and
$\int_{0}^{T}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}$ $\big)^{\top}\mathrm{d}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)=-\int_{0}^{T}\Big(\big({(\delta X)}_{t}^{h}\big)^{\top}A_{t}^{\top}Y^{\alpha,m,\xi}_{t}\Big)\mathrm{d}t$
$-\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}\big((Q_{t}+\bar{Q}_{t})X_{t}^{\alpha,m,\xi}-\bar{Q}_{t}S_{t}m_{t}\big)\mathrm{d}t+\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}Z^{\alpha,m,\xi}_{t}\mathrm{d}W_{t}$
respectively. By (12) $\int_{0}^{\cdot}\big({(\delta X)}_{t}^{h}\big)^{\top}Z^{\alpha,m,\xi}_{t}\mathrm{d}W_{t}$ is a true martingale, hence
$(Y^{\alpha,\boldsymbol{\eta},m}_{T})^{\top}{(\delta X)}_{T}^{h}$ $=-\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}\big((Q_{t}+\bar{Q}_{t})X_{t}^{\alpha,m,\xi}-\bar{Q}_{t}S_{t}m_{t}\big)\mathrm{d}t$
$+\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}B_{t}h_{t}\Big)\mathrm{d}t+M_{T}.$
Taking expectations (the martingale term $M_{T}$ vanishes) gives
$\mathbb{E}\Big[(Y^{\alpha,m,\xi}_{T})^{\top}{(\delta X)}_{T}^{h}\Big]=-\mathbb{E}\!\left[\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}\big((Q_{t}+\bar{Q}_{t})X_{t}^{\alpha,m,\xi}-\bar{Q}_{t}S_{t}m_{t}\big)\mathrm{d}t\right]$
$+\mathbb{E}\!\left[\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}B_{t}h_{t}\Big)\mathrm{d}t\right].$
Using $Y^{\alpha,m,\xi}_{T}=(Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}$, we obtain
$\mathbb{E}\!\Big[\big({(\delta X)}_{T}^{h}\big)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)+\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}\big((Q_{t}+\bar{Q}_{t})X_{t}^{\alpha,m,\xi}-\bar{Q}_{t}S_{t}m_{t}\big)\mathrm{d}t\Big]$
$=\mathbb{E}\!\left[\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}B_{t}h_{t}\Big)\mathrm{d}t\right].$ (14)
Substituting (14) into (13) we obtain
$(\delta\mathcal{J})^{m,\xi,h}(\alpha)=\mathbb{E}\!\left[\int_{0}^{T}\big(B_{t}^{\top}Y^{\alpha,m,\xi}_{t}+R_{t}\alpha_{t}\big)^{\top}h_{t}\,\mathrm{d}t\right].$ (15)
Since $h$ is arbitrary in $\widehat{\mathcal{A}}$, the necessary condition for optimality is
$(\delta\mathcal{J})^{m,\xi,h}(\alpha)=0,\quad\forall\,h\in\widehat{\mathcal{A}}\quad\Longleftrightarrow\quad B_{t}^{\top}Y^{\alpha,m,\xi}_{t}+R_{t}\alpha_{t}=0\quad\text{a.s., for a.e. }t\in[0,T].$
Therefore,
$\alpha_{t}^{*}=-R_{t}^{-1}B_{t}^{\top}Y^{\alpha,m,\xi}_{t}.$ (16)
The stochastic maximum principle thus yields the coupled forward–backward stochastic Volterra system of differential equations
$\left\{\begin{aligned} &{X}^{m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m,\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$. (17)
Remark 3.5.
One can view the search for a solution of (17) as a consistency condition for the control $\alpha$. Indeed, for fixed $\alpha\in\widehat{\mathcal{A}}$, $m\in C([0,T];\mathbb{R}^{d})$, $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d}),$ we have the partially coupled (in the sense that the forward component is independent of the backward) Volterra FBSDE
$\left\{\begin{aligned} &{X}^{\alpha,m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{\alpha,m,\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{\alpha,m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
For the latter we immediately know, by solving first the forward equation and then the backward component, that there exists a unique solution $({X}^{\alpha,m,\xi},{Y}^{\alpha,m,\xi},{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$. Hence we can define the map
$^{*}:\widehat{\mathcal{A}}\rightarrow\widehat{\mathcal{A}},\hskip 5.69046pt\text{where}\hskip 5.69046pt\alpha^{*}_{t}:=-R_{t}^{-1}B_{t}^{\top}Y^{\alpha,m,\xi}_{t},$
and see our problem as searching for fixed points of the map .
System (17), derived by (16) as explained in 3.5, a priori describes a necessary condition for $J^{m,\xi}(\cdot)$ to have a local minimizer in $\widehat{\mathcal{A}}$. Below we show that in fact it provides a necessary and sufficient condition for $J^{m,\xi}(\cdot)$ to have a unique global minimizer in $\widehat{\mathcal{A}}$, as in the non-Volterra setting.
Lemma 3.6 (Optimality of $\alpha^{*}$).
Let $m\in C([0,T];\mathbb{R}^{d})$, $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d})$ and assume that the Volterra system (17) has a solution $(X^{m,\xi},Y^{m,\xi},Z^{m,\xi})$ on $[0,T]$. Then, for every $h\in\widehat{\mathcal{A}}$ we have
$\mathcal{J}^{m,\xi}(\alpha^{*})+\lambda\|h\|_{\mathbb{H}^{2}}^{2}\leq\mathcal{J}^{m,\xi}(\alpha^{*}+h).$
Proof.
As we have already seen, by (8), (9) and (10), for every $h\in\widehat{\mathcal{A}}$ we have
$\mathcal{J}^{m,\xi}(\alpha^{*}+h)-\mathcal{J}^{m,\xi}(\alpha^{*})=R^{m}_{1}({X}^{m,\xi},{(\delta X)}^{h},\alpha^{*},h)+R^{m}_{2}({(\delta X)}^{h},h),$
and because, by (15) and (16) we have by design $R^{m}_{1}({X}^{m,\xi},{(\delta X)}^{h},\alpha^{*},h)=0$, we actually have
$\mathcal{J}^{m,\xi}(\alpha^{*}+h)-\mathcal{J}^{m,\xi}(\alpha^{*})=R^{m}_{2}({(\delta X)}^{h},h).$
So, by (S2), $R_{t}\succ\lambda\text{Id}_{d\times d}$ and $Q_{t},\bar{Q}_{t},Q,\bar{Q}\succeq 0$, we get
$\mathbb{E}\Bigg[\int_{0}^{T}\frac{1}{2}\Big[H_{t}^{\top}(Q_{t}+\bar{Q}_{t})H_{t}\Big]\mathrm{d}t+\frac{1}{2}H_{T}^{\top}(Q+\bar{Q})H_{T}\Bigg]\geq 0$
and
$\lambda\|h\|_{\mathbb{H}^{2}}^{2}\leq\mathbb{E}\Bigg[\int_{0}^{T}h_{t}^{\top}R_{t}h_{t}\mathrm{d}t\Bigg].$
From the above we deduce that for every $H\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$ and $h\in\widehat{\mathcal{A}}$ we have
$\lambda\|h\|_{\mathbb{H}^{2}}^{2}\leq R^{m}_{2}(H,h)$ (18)
and thus $\mathcal{J}^{m}(\alpha^{*})+\lambda\|h\|_{\mathbb{H}^{2}}^{2}\leq\mathcal{J}^{m}(\alpha^{*}+h).$ The proof is complete. ∎
Thus, modulo well-posedness of the Volterra FBSDE (17), we have showed following:
Theorem 3.7.
Let (S1), (S2) and (S3) hold true. Then, for every fixed $m\in C([0,T];\mathbb{R}^{d})$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d})$, the Volterra FBSDE (17) and the linear-quadratic stochastic optimal control problem described by (6), (7) have a unique solution $(\alpha^{*},X^{\alpha^{*},m,\xi})$ [^5] We remind that $X^{m,\xi}=X^{\alpha^{*},m,\xi}$.. Moreover, the optimal control is explicitly given by the formula
$\alpha_{t}^{*}=-R_{t}^{-1}B_{t}^{\top}Y^{m,\xi}_{t},$
where $Y^{m,\xi}$ is the solution of the backward component of (17).
Proof.
Immediate from 3.2, 3.4, Lemma 3.6. Well-posedness of (17) is established in Theorem 3.9 below. ∎
3.3 Well posedness of the Volterra FBSDE
In this subsection we complete the proof of Theorem 3.7 by establishing existence and uniqueness of the Volterra FBSDE (17) under (S1), (S2) and (S3).
To this end, we first introduce the processes:
$\widehat{X}^{m,\xi}_{t}:=\mathcal{K}(0,t)\big({X}^{m,\xi}_{t}-M(m)_{t}\big),\hskip 2.84544pt\widehat{Y}^{m,\xi}_{t}:=\mathcal{K}(t,0)^{\top}{Y}^{m,\xi}_{t},$ (19)
$\widehat{\Sigma}_{t}:=\mathcal{K}(0,t)\Sigma_{t},\hskip 2.84544pt\widehat{Z}_{t}^{m,\xi}:=\mathcal{K}(s,0)^{\top}Z_{s}.$ (20)
In view of S3, the latter transform (17) into the following standard FBSDE with affine structure:
$\left\{\begin{aligned} &\widehat{X}^{m,\xi}_{t}=\xi+\int_{0}^{t}\mathcal{K}(0,s)\big(A_{s}\mathcal{K}(s,0)\widehat{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}\mathcal{K}(0,s)^{\top}\widehat{Y}^{m,\xi}_{s}+A_{s}M(m)_{s}+C_{s}m_{s}\big)\,\mathrm{d}s\\ &\quad\quad\quad\quad+\int_{0}^{t}\widehat{\Sigma}_{s}\mathrm{d}W_{s},\\ &\widehat{Y}^{m,\xi}_{t}=\mathcal{K}(T,0)^{\top}\big((Q+\bar{Q})\mathcal{K}(T,0)\widehat{X}_{T}^{m,\xi}+(Q+\bar{Q})M(m)_{T}-\bar{Q}Sm_{T}\big)\\ &\quad+\int_{t}^{T}\mathcal{K}(s,0)^{\top}\Big((Q_{s}+\bar{Q}_{s})\mathcal{K}(s,0)\widehat{X}_{s}^{m,\xi}+(Q_{s}+\bar{Q}_{s})M(m)_{s}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathcal{K}(0,s)^{\top}\widehat{Y}^{m,\xi}_{s}\Big)\mathrm{d}s\\ &\quad\quad-\int_{t}^{T}\widehat{Z}_{s}^{m,\xi}\mathrm{d}W_{s}.\end{aligned}\right$. (21)
Riccati approach
In order to solve (21) we are going to assume a decoupling field $V$ with affine structure. Subsequently, we verify that the results we obtain indeed satisfy (21). The ansatz is
$V(t,x):=P_{t}x+\Pi_{t},\hskip 2.84544pt\forall(t,x)\in[0,T]\times\mathbb{R}^{d}$
and
$\widehat{Y}^{m,\xi}_{t}=P_{t}\widehat{X}^{m,\xi}_{t}+\Pi_{t}.$ (22)
Of course, we have so far only postulated the form of $V.$ In order to properly define it, we must specify the coefficients $P:[0,T]\rightarrow\mathbb{R}^{d\times d},\Pi:[0,T]\rightarrow\mathbb{R}^{d}$. By taking expectations on (21) and (22) we get
$\left\{\begin{aligned} &\mathbb{E}[\widehat{X}^{m,\xi}_{t}]=\mathbb{E}[\xi]+\int_{0}^{t}\mathcal{K}(0,s)\big(A_{s}\mathcal{K}(s,0)\mathbb{E}[\widehat{X}^{m,\xi}_{s}]-B_{s}R_{s}^{-1}B_{s}^{\top}\mathcal{K}(0,s)^{\top}\mathbb{E}[\widehat{Y}^{m,\xi}_{s}]\\ &\quad\quad\quad\quad+A_{s}M(m)_{s}+C_{s}m_{s}\big)\,\mathrm{d}s\\ &\mathbb{E}[\widehat{Y}^{m,\xi}_{t}]=\mathcal{K}(T,0)^{\top}\big((Q+\bar{Q})\mathcal{K}(T,0)\mathbb{E}[\widehat{X}_{T}^{m,\xi}]+(Q+\bar{Q})M(m)_{T}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,0)^{\top}\Big((Q_{s}+\bar{Q}_{s})\mathcal{K}(s,0)\mathbb{E}[\widehat{X}_{s}^{m,\xi}]+(Q_{s}+\bar{Q}_{s})M(m)_{s}\\ &\quad\quad\quad\quad-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathcal{K}(0,s)^{\top}\mathbb{E}[\widehat{Y}^{m,\xi}_{s}]\Big)\mathrm{d}s.\end{aligned}\right$. (23)
and
$\mathbb{E}[\widehat{Y}^{m,\xi}_{t}]=P_{t}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]+\Pi_{t}.$ (24)
For ease of notation we define, for every $t\in[0,T],$ the transformed coefficients
$\widehat{A}_{t}=\mathcal{K}(0,t)A_{t}\mathcal{K}(t,0),\hskip 2.84544pt\widehat{B}_{t}=\mathcal{K}(0,t)B_{t}\mathcal{K}(t,0),\hskip 2.84544pt\widehat{R}_{t}=\mathcal{K}(t,0)^{\top}R_{t}\mathcal{K}(t,0),\hskip 2.84544pt\widehat{Q}_{t}=\mathcal{K}(t,0)^{\top}Q_{t}\mathcal{K}(t,0),$ (25)
$\widehat{\bar{Q}}_{t}=\mathcal{K}(t,0)^{\top}\bar{Q}_{t}\mathcal{K}(t,0),\hskip 2.84544pt\widehat{Q}=\mathcal{K}(T,0)^{\top}Q\mathcal{K}(T,0),\hskip 2.84544pt\widehat{\bar{Q}}=\mathcal{K}(T,0)^{\top}\bar{Q}\mathcal{K}(T,0),$
$\widehat{D}^{m}_{t}=\mathcal{K}(0,t)A_{t}M(m)_{t}+\mathcal{K}(0,t)C_{t}m_{t}.$
With the help of (25) we rewrite (23) as
$\left\{\begin{aligned} &\mathbb{E}[\widehat{X}^{m,\xi}_{t}]=\mathbb{E}[\xi]+\int_{0}^{t}\big(\widehat{A}_{s}\mathbb{E}[\widehat{X}^{m,\xi}_{s}]-\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}\mathbb{E}[\widehat{Y}^{m,\xi}_{s}]+\widehat{D}^{m}_{s}\big)\,\mathrm{d}s\\ &\mathbb{E}[\widehat{Y}^{m,\xi}_{t}]=(\widehat{Q}+\widehat{\bar{Q}})\mathbb{E}[\widehat{X}_{T}^{m,\xi}]+(Q+\bar{Q})M(m)_{T}-\widehat{\bar{Q}}\mathcal{K}(0,T)Sm_{T}\\ &\quad\quad\quad\quad+\int_{t}^{T}\Big((\widehat{Q}_{s}+\widehat{\bar{Q}}_{s})\mathbb{E}[\widehat{X}_{s}^{m,\xi}]+(\widehat{A}_{s})^{\top}\mathbb{E}[\widehat{Y}^{m,\xi}_{s}]\\ &\quad\quad\quad\quad+(\widehat{Q}_{s}+\widehat{\bar{Q}}_{s})\mathcal{K}(0,s)M(m)_{s}-\widehat{\bar{Q}}_{s}\mathcal{K}(0,s)S_{s}m_{s}\Big)\mathrm{d}s.\end{aligned}\right$. (26)
Remark 3.8.
It is immediate that $\widehat{Q},\widehat{\bar{Q}},\widehat{Q}_{t},\widehat{\bar{Q}}_{t}\succeq 0$. For $\widehat{R}_{t}$ we have $\widehat{R}_{t}\succ\lambda\big(\mathcal{K}(t,0)^{\top}\mathcal{K}(t,0)\big)$. Because $\mathcal{K}(t,0)$ is invertible and continuous over $[0,T]$, which is compact, there exists $C_{\mathcal{K},T}>0$ such that for every $x\in\mathbb{R}^{d}$ we have $C_{\mathcal{K},T}|x|\leq|\mathcal{K}(t,0)x|$. Hence, $\widehat{R}_{t}\succ(\lambda C_{\mathcal{K},T})\text{Id}_{d\times d}$.
Now, by differentiating (24) and using the forward part of (26), together with (24), we get
$\dot{\mathbb{E}[\widehat{Y}^{m,\xi}_{t}]}=\dot{P_{t}}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]+P_{t}\dot{\mathbb{E}[\widehat{X}^{m,\xi}_{t}]}+\dot{\Pi}_{t}$
$\quad=\dot{P_{t}}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]+P_{t}\big(\widehat{A}_{t}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]-\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}\mathbb{E}[\widehat{Y}^{m,\xi}_{t}]+\widehat{D}^{m}_{t}\big)+\dot{\Pi}_{t}$
$\quad=\dot{P_{t}}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]+P_{t}\big(\widehat{A}_{t}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]-\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}\big(P_{t}\mathbb{E}[\widehat{X}^{m,\xi}_{t}]+\Pi_{t}\big)+\widehat{D}^{m}_{t}\big)+\dot{\Pi}_{t}$
$\quad=\Big[\dot{P_{t}}+P_{t}\widehat{A}_{t}-P_{t}\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}P_{t}\Big]\mathbb{E}[\widehat{X}^{m,\xi}_{t}]+\Big[\dot{\Pi}_{t}-\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}\Pi_{t}+P_{t}\Pi_{t}+P_{t}\widehat{D}^{m}_{t}\Big].$
And by the backward part of (26), together with (24), we get
$\dot{\mathbb{E}[\widehat{Y}^{m,\xi}_{t}]}=-\Big[(\widehat{Q}_{t}+\widehat{\bar{Q}}_{t})+(\widehat{A}_{t})^{\top}P_{t}\Big]\mathbb{E}[\widehat{X}_{t}^{m,\xi}]-\Big[(\widehat{A}_{t})^{\top}\Pi_{t}-\widehat{\bar{Q}}_{t}\mathcal{K}(0,t)S_{t}m_{t}\Big],$
hence
$0=\Big[\dot{P_{t}}+P_{t}\widehat{A}_{t}+(\widehat{A}_{t})^{\top}P_{t}-P_{t}\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}P_{t}+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{t})\Big]\mathbb{E}[\widehat{X}^{m,\xi}_{t}]$
$\hskip 2.84544pt+\Big[\dot{\Pi}_{t}+\big(P_{t}+(\widehat{A}_{t})^{\top}-\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}\big)\Pi_{t}+P_{t}\widehat{D}^{m}_{t}+(\widehat{Q}_{s}+\widehat{\bar{Q}}_{t})\mathcal{K}(0,t)M(m)_{t}-\widehat{\bar{Q}}_{t}\mathcal{K}(0,t)S_{t}m_{t}\Big].$
From the above equation, and because $\mathbb{E}[\xi]$ may take any value of $\mathbb{R}^{d}$ at (26), we derive the system of backward matrix Ordinary Differential Equations
$\left\{\begin{aligned} &\dot{P_{t}}+P_{t}\widehat{A}_{t}+(\widehat{A}_{t})^{\top}P_{t}-P_{t}\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}P_{t}+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{t})=0,\\ &\dot{\Pi}_{t}+\big(P_{t}+(\widehat{A}_{t})^{\top}-\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}\big)\Pi_{t}+P_{t}\widehat{D}^{m}_{t}+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{t})\mathcal{K}(0,t)M(m)_{t}-\widehat{\bar{Q}}_{t}\mathcal{K}(0,t)S_{t}m_{t}=0,\end{aligned}\right$. (27)
with terminal conditions $P_{T}=\widehat{Q}+\widehat{\bar{Q}}$ and $\Pi_{T}=(Q+\bar{Q})M(m)_{T}-\widehat{\bar{Q}}\mathcal{K}(0,T)Sm_{T}$.
The first equation of the above system is a symmetric (see 3.8) Riccati matrix differential equation with terminal condition, and the second is a non-homogeneous linear matrix differential equation with terminal condition.
Assuming that the Riccati equation has a unique solution on $[0,T]$, then the second equation for $\Pi_{t}$ is obviously also uniquely solvable on $[0,T]$, and the decoupling field $V(t,x)$ is properly defined. Next, using (22) we write the forward equation of (21) as
$\widehat{X}^{m,\xi}_{t}=\xi+\int_{0}^{t}$ $\mathcal{K}(0,s)\big(A_{s}\mathcal{K}(s,0)\widehat{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}\mathcal{K}(0,s)^{\top}(P_{s}\widehat{X}^{m,\xi}_{s}+\Pi_{s})\,\mathrm{d}s$
$+\int_{0}^{t}\mathcal{K}(0,s)\big(A_{s}M(m)_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\widehat{\Sigma}_{s}\mathrm{d}W_{s},$
which is an affine equation with respect to $\widehat{X}^{m,\xi}_{t}$ that obviously has a unique solution $\widehat{X}^{m,\xi}_{t}\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$. Finally, one can directly verify via Itô’s formula (see for example [50, Corollary 6.4] ($V(t,x)\in C^{1,2}([0,T]\times\mathbb{R}^{d};\mathbb{R}^{d})$), and the terminal conditions for $P_{T},\Pi_{T}$, that $V(t,\widehat{X}^{m,\xi}_{t})$ satisfies the backward equation of (21). In other words, we have found a solution of (21) on the whole $[0,T]$, where the backward component is expressed via an affine decoupling field with the help of the forward component. Hence, to get uniqueness for the solution of (21) we apply [19, Proposition 4.8], and so we have arrived at the following.
Theorem 3.9.
Let (S1), (S2) and (S3) hold true, then for every fixed $m\in C([0,T];\mathbb{R}^{d})$ and $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d})$, the Volterra FBSDE (17) has a unique solution $({X}^{m,\xi},{Y}^{m,\xi},{Z}^{m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$.
Proof.
What remains to be proved is that the symmetric Riccati equation
$\dot{P_{t}}+P_{t}\widehat{A}_{t}+(\widehat{A}_{t})^{\top}P_{t}-P_{t}\widehat{B}_{t}(\widehat{R}_{t})^{-1}(\widehat{B}_{t})^{\top}P_{t}+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{t})=0,\hskip 5.69046ptP_{T}=\widehat{Q}+\widehat{\bar{Q}}$
has a unique global solution on $[0,T]$. But this follows [^6] We remind again 3.8. from A.2 with the rough components equal to zero. ∎
4 Solving the rough LQ MFG
4.1 Equivalence of Volterra fixed point step with Volterra FBODE
Having dealt with the optimal control phase of the problem in Section 3, this section is devoted to the fixed point step. We will show that, as in the classical case, existence and uniqueness of the fixed point, is equivalent to existence and uniqueness of a Volterra forward-backward ODE (FBODE) system.
We fix an initial condition $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0};\mathbb{R}^{d}).$For ease of notation, we are going to omit the dependence on $\xi,$ whenever possible, in this section. Additionally, we remind that (S1), (S2) and (S3) remain valid.
From 3.7, for every $m\in C([0,T];\mathbb{R}^{d})$ the LQ stochastic Volterra optimal control problem has a unique solution, and the optimal state satisfies, along with its adjoint state, the Volterra system
$\left\{\begin{aligned} &{X}^{m}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{m}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m}-\bar{Q}Sm_{T}\big)\\ &\quad\quad\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
Taking expectations on the above system we have
$\left\{\begin{aligned} &\mathbb{E}[{X}^{m}_{t}]=\mathcal{K}(t,0)\mathbb{E}[\xi]+M(m)_{t}+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}\mathbb{E}[{X}^{m}_{s}]-B_{s}R_{s}^{-1}B_{s}^{\top}\mathbb{E}[Y^{m}_{s}]+C_{s}m_{s}\big)\,\mathrm{d}s,\\ &\mathbb{E}[Y^{m}_{t}]=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})\mathbb{E}[X_{T}^{m}]-\bar{Q}Sm_{T}\big)\\ &\quad\quad\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})\mathbb{E}[X_{s}^{m}]-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathbb{E}[Y^{m}_{s}]\Big)\mathrm{d}s.\end{aligned}\right$.
If $m^{0}$ is a fixed point of the map $F:C([0,T];\mathbb{R}^{d})\rightarrow C([0,T];\mathbb{R}^{d})$, where
$F(m)_{t}:=\mathbb{E}[{X}^{m}_{t}],$ (28)
then it satisfies the consistency equation $m_{t}=\mathbb{E}[{X}^{m}_{t}]$ for all $t\in[0,T].$ Plugging this equation to the system above, we see that $\bar{x}_{t}:=\mathbb{E}[{X}^{m^{0}}_{t}],\bar{y}_{t}:=\mathbb{E}[{Y}^{m^{0}}_{t}]$ satisfy
$\left\{\begin{aligned} &\bar{x}_{t}=\mathcal{K}(t,0)\mathbb{E}[\xi]+M(\bar{x})_{t}+\int_{0}^{t}\mathcal{K}(t,s)\big((A_{s}+C_{s})\bar{x}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}\bar{y}_{s}\big)\,\mathrm{d}s,\\ &\bar{y}_{t}=\mathcal{K}(T,t)^{\top}(Q+\bar{Q}-\bar{Q}S)\bar{x}_{T}+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s}-\bar{Q}_{s}S_{s})\bar{x}_{s}+A_{s}^{\top}\bar{y}_{s}\Big)\mathrm{d}s.\end{aligned}\right$. (29)
Remark 4.1.
The term $M(\bar{x})_{t}$ in the forward component of system (29) makes it more difficult to solve it compared to the classical case. Indeed, one cannot apply the same transformation as in the beginning of Section 3.3 above.
Nevertheless, (29) completely characterizes the existence and uniqueness in terms of the fixed points of $F$.
Theorem 4.2.
There exists a one-to-one correspondence between the fixed points of map $F$ (28) and the solutions of system (29). Specifically, every fixed point is the forward component of a solution of (29), and vise versa, every forward component of a solution of (29) is a fixed point of $F$.
Proof.
First we assume that $F$ has a fixed point $m^{0}$. From the analysis before 4.1 we have
$\begin{cases}\mathbb{E}[{X}^{m^{0}}_{t}]=\mathcal{K}(t,0)\mathbb{E}[\xi]+M(\mathbb{E}[{X}^{m^{0}}_{\cdot}])_{t}+\int_{0}^{t}\mathcal{K}(t,s)\big((A_{s}+C_{s})\mathbb{E}[{X}^{m^{0}}_{s}]-B_{s}R_{s}^{-1}B_{s}^{\top}\mathbb{E}[Y^{m^{0}}_{s}]\big)\,\mathrm{d}s,\\[3.00003pt] \mathbb{E}[Y^{m^{0}}_{t}]=\mathcal{K}(T,t)^{\top}(Q+\bar{Q}-\bar{Q}S)\mathbb{E}[X_{T}^{m^{0}}]\\ \hskip 119.50148pt+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s}-\bar{Q}_{s}S_{s})\mathbb{E}[X_{s}^{m^{0}}]+A_{s}^{\top}\mathbb{E}[Y^{m^{0}}_{s}]\Big)\mathrm{d}s,\end{cases}$
and thus (29) is solved by the pair $\big(\mathbb{E}[{X}^{m^{0}}_{\cdot}],\mathbb{E}[{Y}^{m^{0}}_{\cdot}]\big)=\big(m^{0},\mathbb{E}[{Y}^{m^{0}}_{\cdot}]\big)$. In other words, we have a mapping $m^{0}\mapsto\big(m^{0},\mathbb{E}[{Y}^{m^{0}}_{\cdot}]\big)$ from fixed points to solution of the FBODE.
For the opposite direction, let $(\bar{x}^{0},\bar{y}^{0})$ be a solution of (29). From 3.9 there exists a unique triplet $(X^{\bar{x}^{0}},Y^{\bar{x}^{0}},Z^{\bar{x}^{0}})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$ such that
$\begin{cases}{X}^{\bar{x}^{0}}_{t}=\mathcal{K}(t,0)\xi+M(\bar{x}^{0})_{t}\\ \hskip 79.6678pt+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{\bar{x}^{0}}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{\bar{x}^{0}}_{s}+C_{s}\bar{x}^{0}_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\[3.00003pt] Y^{\bar{x}^{0}}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{\bar{x}^{0}}-\bar{Q}S\bar{x}^{0}_{T}\big)\\ \hskip 51.21504pt+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\bar{x}^{0}}-\bar{Q}_{s}S_{s}\bar{x}^{0}_{s}+A_{s}^{\top}Y^{\bar{x}^{0}}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{\bar{x}^{0}}_{s}\mathrm{d}W_{s}.\end{cases}$
We want to show that $F(\bar{x}^{0})_{t}=\mathbb{E}[{X}^{\bar{x}^{0}}_{t}]=\bar{x}^{0}_{t}.$ By taking expectations above we have that the pair $\big(\mathbb{E}[{X}^{\bar{x}^{0}}_{\cdot}],\mathbb{E}[{Y}^{\bar{x}^{0}}_{\cdot}]\big)$ satisfies
$\begin{cases}\mathbb{E}[{X}^{\bar{x}^{0}}_{t}]=\mathcal{K}(t,0)\mathbb{E}[\xi]+M(\bar{x}^{0})_{t}+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}\mathbb{E}[{X}^{\bar{x}^{0}}_{s}]-B_{s}R_{s}^{-1}B_{s}^{\top}\mathbb{E}[Y^{\bar{x}^{0}}_{s}]+C_{s}\bar{x}^{0}_{s}\big)\,\mathrm{d}s,\\[3.00003pt] \mathbb{E}[Y^{\bar{x}^{0}}_{t}]=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})\mathbb{E}[X_{T}^{\bar{x}^{0}}]-\bar{Q}S\bar{x}^{0}_{T}\big)\\ \hskip 51.21504pt+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})\mathbb{E}[X_{s}^{\bar{x}^{0}}]-\bar{Q}_{s}S_{s}\bar{x}^{0}_{s}+A_{s}^{\top}\mathbb{E}[Y^{\bar{x}^{0}}_{s}]\Big)\mathrm{d}s.\end{cases}$
Therefore, the triplet $\big(\mathbb{E}[{X}^{\bar{x}^{0}}_{\cdot}]-\bar{x}^{0},\mathbb{E}[{Y}^{\bar{x}^{0}}_{\cdot}]-\bar{y}^{0},0\big)$ satisfies
$\begin{cases}\mathbb{E}[{X}^{\bar{x}^{0}}_{t}]-\bar{x}^{0}_{t}=\int_{0}^{t}\mathcal{K}(t,s)\Big(A_{s}\big(\mathbb{E}[{X}^{\bar{x}^{0}}_{s}]-\bar{x}^{0}_{s}\big)-B_{s}R_{s}^{-1}B_{s}^{\top}\big(\mathbb{E}[Y^{\bar{x}^{0}}_{s}]-\bar{y}^{0}_{s}\big)\Big)\,\mathrm{d}s,\\[3.00003pt] \mathbb{E}[Y^{\bar{x}^{0}}_{t}]-\bar{y}^{0}_{t}=\mathcal{K}(T,t)^{\top}(Q+\bar{Q})\big(\mathbb{E}[X_{T}^{\bar{x}^{0}}]-\bar{x}^{0}_{T}\big)\\ \hskip 51.21504pt+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})\big(\mathbb{E}[X_{s}^{\bar{x}^{0}}]-\bar{x}^{0}_{s}\big)+A_{s}^{\top}\big(\mathbb{E}[Y^{\bar{x}^{0}}_{s}]-\bar{y}^{0}_{s}\big)\Big)\mathrm{d}s.\end{cases}$
This last system has the form of (17) with $\xi=0,M(\cdot)=0,\Sigma_{t}=0,C_{t}=0,S_{t}=0,S=0.$ In view of 3.9 it has a unique solution which is equal to $(0,0,0)$. Thus, $\bar{x}^{0}$ is a fixed point of F and $(\bar{x}^{0},\bar{y}^{0})=\big(\bar{x}^{0},\mathbb{E}[{Y}^{\bar{x}^{0}}_{\cdot}]\big)$. ∎
4.2 Connecting the Volterra and rough frameworks
In this section we establish the equivalence between the Volterra and the rough settings. For this purpose we are going to restrict the space of admissible controls to
$\mathcal{A}=\bigg\{\alpha\in\widehat{\mathcal{A}}:\sup_{t\in[0,T]}\mathbb{E}[|\alpha_{t}|^{k}]<\infty,\hskip 2.84544pt\text{for every}\hskip 2.84544ptk\in[2,\infty)\bigg\},$ (30)
while we also assume that $\Sigma\in\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$ is such that $\sup_{t\in[0,T]}\mathbb{E}[|\Sigma_{t}|^{k}]<\infty$ for every $k\in[2,\infty)$.
Let $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p}),p\in\mathbb{N}$, with $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. Recall from Theorem 2.9 that the coefficients $A^{1},C^{1}$ in the state dynamics are taken to be controlled rough paths. Thus we impose:
  1. (S1-R)
    There exists $\big({A}^{1},(A^{1})^{\prime}\big),\big({C}^{1},(C^{1})^{\prime}\big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$.
Turning to the choice of the Volterra kernel, (S1-R) along with 2.9 guarantee that, for any $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$, the deterministic controlled rough paths $\big(U^{\boldsymbol{\eta}},(U^{\boldsymbol{\eta}})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ and $\big(M^{\boldsymbol{\eta},(m,m^{\prime})},(M^{\boldsymbol{\eta},(m,m^{\prime})})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ respectively solve the linear RDEs
$U^{\boldsymbol{\eta}}_{t\leftarrow s}$ $=\text{Id}_{d\times d}+\int_{s}^{t}A^{1}_{r}U^{\boldsymbol{\eta}}_{r\leftarrow s}\mathrm{d}\boldsymbol{\eta}_{r},$ (31)
$M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$ $=\int_{0}^{t}\big(A^{1}_{r}M^{\boldsymbol{\eta},(m,m^{\prime})}_{r}+C^{1}_{r}m_{r}\big)\mathrm{d}\boldsymbol{\eta}_{r}.$
The matrix $U^{\boldsymbol{\eta}}_{t\leftarrow s}$ is invertible for every $t\in[0,T]$ (this follows e.g. by applying A.3). Its inverse $(U^{\boldsymbol{\eta}}_{t\leftarrow s})^{-1}=:U^{\boldsymbol{\eta}}_{s\leftarrow t}$, where $\big(U^{\boldsymbol{\eta}}_{s\leftarrow t},(U^{\boldsymbol{\eta}}_{s\leftarrow t})^{\prime}\big)\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ is the solution of
$U^{\boldsymbol{\eta}}_{s\leftarrow t}=\text{Id}_{d\times d}-\int_{s}^{t}U^{\boldsymbol{\eta}}_{s\leftarrow r}A^{1}_{r}\mathrm{d}\boldsymbol{\eta}_{r}.$
Remark 4.3.
Note that, for every $s,t\in[0,T]$ with $s\leq t$ we have
$(U^{\boldsymbol{\eta}}_{t\leftarrow s})^{\top}$ $=\text{Id}_{d\times d}+\int_{s}^{t}(U^{\boldsymbol{\eta}}_{r\leftarrow s})^{\top}(A^{1}_{r})^{\top}\mathrm{d}\boldsymbol{\eta}_{r},$
$(U^{\boldsymbol{\eta}}_{s\leftarrow t})^{\top}$ $=\text{Id}_{d\times d}-\int_{s}^{t}(A^{1}_{r})^{\top}(U^{\boldsymbol{\eta}}_{s\leftarrow r})^{\top}\mathrm{d}\boldsymbol{\eta}_{r}.$
We shall need the following instance of Duhamel’s principle. From A.3, applied in ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$, we get that
${U}^{\boldsymbol{\eta}}_{0\leftarrow t}M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}=\int_{0}^{t}U^{\boldsymbol{\eta}}_{0\leftarrow s}C^{1}_{s}m_{s}\mathrm{d}\boldsymbol{\eta}_{s}.$ (32)
Now, let us denote with ${{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ the unique solution of (1), see 2.9, and with $\widetilde{{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}$ the unique solution, see 3.2, of equation (6) where $M(m)_{t}$ is replaced by $M^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$ and
$\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(t,s):={U}^{\boldsymbol{\eta}}_{t\leftarrow 0}{U}^{\boldsymbol{\eta}}_{0\leftarrow s},$ (33)
for every $t,s\in[0,T]$. Since $(A^{1},(A^{1})^{\prime})$ is fixed throughout this section, we shall omit it from our notation and write $\mathcal{K}^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime})}(\cdot,\cdot)$ instead of $\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot)$.
Remark 4.4.
The Volterra kernel $\mathcal{K}^{\boldsymbol{\eta}}$ satisfies defined above satisfies Assumption S1. This can be verified directly from the flow and continuity properties of the linear RDE 31.
We are now ready to state the desired equivalence result for the state processes via Duhamel’s principle/variation of constants.
Theorem 4.5.
Let (S1-R), (S2) hold true, and with the above clarifications, for every $t\in[0,T]$, we have
$\mathbb{P}\big[{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}=\widetilde{{X}}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}\big]=1.$
Proof.
Direct application of the rough product formula, A.3, to the product ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}{X}^{\boldsymbol{\eta},\alpha,(m,m^{\prime}),\xi}_{t}$ with ${U}^{\boldsymbol{\eta}}_{0\leftarrow t}$ satisfying (32).
Definition 4.6.
Under (S1-R), for every $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$, we have (see Lemma 2.6)
$\big(Z^{\boldsymbol{\eta},(m,m^{\prime})},(Z^{\prime})^{\boldsymbol{\eta},(m,m^{\prime})}\big):=$
$\bigg(\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)C^{1}m\;,\;\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)^{\prime}C^{1}m+\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)(C^{1})^{\prime}m+\mathcal{K}^{\boldsymbol{\eta}}(0,\cdot)C^{1}m^{\prime}\bigg)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{p\times d}\big).$ (34)
In turn, we define
$M^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime}),(C^{1},(C^{1})^{\prime})}\big((m,m^{\prime})\big)_{t}:=\mathcal{K}^{\boldsymbol{\eta}}(t,0)\int_{0}^{t}\big(Z^{\boldsymbol{\eta},(m,m^{\prime})},(Z^{\prime})^{\boldsymbol{\eta},(m,m^{\prime})}\big)_{s}\mathrm{d}\boldsymbol{\eta}_{s}.$ (35)
Since $(A^{1},(A^{1})^{\prime}),(C^{1},(C^{1})^{\prime})$ are fixed in this section, for ease of notation, we shall denote $M^{\boldsymbol{\eta},(A^{1},(A^{1})^{\prime}),(C^{1},(C^{1})^{\prime})}(\cdot)$ by $M^{\boldsymbol{\eta}}(\cdot)$
Hence, for fixed $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, under (S1-R), (S2) and (S3), from the Volterra formulation where $M(m)$ is replaced from $M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)$, the stochastic optimal control problem has a unique solution. Furthermore, by 3.9 and 4.5, for the optimal state and optimal control $\big(X^{(m,m^{\prime}),\xi},\alpha^{*,(m,m^{\prime}),\xi}\big)$, of the rough stochastic optimal control problem we get the representations
$\left\{\begin{aligned} &X^{(m,m^{\prime}),\xi}_{t}=\mathcal{K}^{\boldsymbol{\eta}}(t,0)\xi+M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{t}\\ &\quad\quad\quad+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(t,s)\big(A_{s}X^{(m,m^{\prime}),\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{(m,m^{\prime}),\xi}_{t}=\mathcal{K}^{\boldsymbol{\eta}}(T,t)^{\top}\big((Q+\bar{Q})X^{(m,m^{\prime}),\xi}_{T}-\bar{Q}Sm_{T}\big)\\ &\quad\quad\quad\quad+\int_{t}^{T}\mathcal{K}^{\boldsymbol{\eta}}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X^{(m,m^{\prime}),\xi}_{s}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}\Big)\mathrm{d}s\\ &\quad\quad\quad\quad\quad-\int_{t}^{T}\mathcal{K}^{\boldsymbol{\eta}}(s,t)^{\top}Z^{(m,m^{\prime}),\xi}_{s}\mathrm{d}W_{s},\\ &\alpha^{*,(m,m^{\prime}),\xi}_{t}=-R_{t}^{-1}B_{t}^{\top}Y^{(m,m^{\prime}),\xi}_{t}.\end{aligned}\right$. (36)
Remark 4.7.
It is important to note that
$\sup_{t\in[0,T]}\mathbb{E}\Big[\big|X^{(m,m^{\prime}),\xi}_{t}\big|^{k}\Big]+\sup_{t\in[0,T]}\mathbb{E}\Big[\big|Y^{(m,m^{\prime}),\xi}_{t}\big|^{k}\Big]+\sup_{t\in[0,T]}\mathbb{E}\Big[\big|\alpha^{*,(m,m^{\prime}),\xi}_{t}\big|^{k}\Big]<\infty,\hskip 5.69046pt\text{for every}\hskip 2.84544ptk\in[2,\infty).$
Indeed, by integrability of $\xi$ and the fact that the transformed problem (recall (19)) is solved through a (deterministic) decoupling field, standard SDE estimates, see e.g. [54, Theorem 3.4.3], imply
$\sup_{t\in[0,T]}\mathbb{E}\Big[\big|\widehat{X}^{(m,m^{\prime}),\xi}_{t}\big|^{k}\Big]+\sup_{t\in[0,T]}\mathbb{E}\Big[\big|\widehat{Y}^{(m,m^{\prime}),\xi}_{t}\big|^{k}\Big]<\infty,\hskip 5.69046pt\text{for every}\hskip 2.84544ptk\in[2,\infty).$
Hence, the desired result follows.
We close this section with the analogue of 3.9.
Theorem 4.8.
Let (S1-R), (S2), (S3) hold true. Then, for every $(m,m^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\xi\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})$, the rough FBSDE
$\left\{\begin{aligned} &X^{(m,m^{\prime}),\xi}_{t}=\xi+\int_{0}^{t}\big(A^{1}_{s}X^{(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad\quad\quad+\int_{0}^{t}\big(A_{s}X^{(m,m^{\prime}),\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{(m,m^{\prime}),\xi}_{t}=(Q+\bar{Q})X^{(m,m^{\prime}),\xi}_{T}-\bar{Q}Sm_{T}+\int_{t}^{T}(A^{1}_{s})^{\top}Y^{(m,m^{\prime}),\xi}_{s}\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad\quad\quad+\int_{t}^{T}\big((Q_{s}+\bar{Q}_{s})X^{(m,m^{\prime}),\xi}_{s}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}\big)\mathrm{d}s-\int_{t}^{T}Z^{(m,m^{\prime}),\xi}_{s}\mathrm{d}W_{s},\end{aligned}\right$. (37)
has a unique solution
$\Big(\big(X^{(m,m^{\prime}),\xi},(X^{(m,m^{\prime}),\xi})^{\prime}\big),\big(Y^{(m,m^{\prime}),\xi},(Y^{(m,m^{\prime}),\xi})^{\prime}\big),Z^{(m,m^{\prime}),\xi}\Big)$
$\in\bigcap_{k=2}^{\infty}\Big(\textbf{D}^{\gamma,\beta}_{\eta}\mathbb{L}^{k}\big([0,T];\mathbb{R}^{d}\big)\times\textbf{D}^{\gamma,\beta}_{\eta}\mathbb{L}^{k}\big([0,T];\mathbb{R}^{d}\big)\times\mathbb{H}^{k}([0,T];\mathbb{R}^{d\times q})\Big).$
with
$\sup_{t\in[0,T]}\mathbb{E}\big[|Z^{(m,m^{\prime}),\xi}_{t}|^{k}\big]<\infty,\hskip 5.69046pt\text{for every}\hskip 2.84544ptk\in[2,\infty).$
Furthermore, the solution (37) almost surely coincides with the one of (36).
Proof.
First we are going to establish existence of such a solution. From the discussion above, and specifically 3.9, there exist unique processes $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},$ $(Z^{1})^{(m,m^{\prime}),\xi}$, with the appropriate integrabilities for $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},$ (see 4.7) that solve (36). We are going to show that $X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi}$ are also part of a solution for (37) along with an appropriate $Z^{(m,m^{\prime}),\xi}$.
Keeping $Y^{(m,m^{\prime}),\xi}$ fixed, we consider the decoupled Volterra and RSDEs, respectively given by
$X^{(m,m^{\prime}),\xi}_{t}=\mathcal{K}^{\boldsymbol{\eta}}(t,0)\xi+M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{t}$
$\hskip 56.9055pt+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(t,s)\big(A_{s}X^{(m,m^{\prime}),\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(t,s)\Sigma_{s}\mathrm{d}W_{s},$
$\widetilde{X}^{(m,m^{\prime}),\xi}_{t}=\xi+\int_{0}^{t}\big(A^{1}_{s}\widetilde{X}^{(m,m^{\prime}),\xi}_{s}+C^{1}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}$
$\hskip 113.81102pt+\int_{0}^{t}\big(A_{s}\widetilde{X}^{(m,m^{\prime}),\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s}.$
From 2.9 both equations admit unique solutions and an application of 4.5 yields $\mathbb{P}\big[X^{(m,m^{\prime}),\xi}=\widetilde{X}^{(m,m^{\prime}),\xi}\big]=1$.
Now let us argue for the backward component. From Section 3.3, there exist $P^{\boldsymbol{\eta}}_{t},\Pi^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$ solutions of (27), recall (25), such that
$Y^{(m,m^{\prime}),\xi}_{t}=\mathcal{K}^{\boldsymbol{\eta}}(0,t)^{\top}P^{\boldsymbol{\eta}}_{t}\mathcal{K}^{\boldsymbol{\eta}}(0,t)\big({X}^{(m,m^{\prime}),\xi}_{t}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{t}\big)+\mathcal{K}^{\boldsymbol{\eta}}(0,t)^{\top}\Pi^{\boldsymbol{\eta},(m,m^{\prime})}_{t}$. (38)
For the rest of the proof we are going to denote $P^{\boldsymbol{\eta}},\Pi^{\boldsymbol{\eta},(m,m^{\prime})}$ by $P,\Pi$ for ease of notation. We will apply A.3 several times in order to see that $Y^{(m,m^{\prime}),\xi}$ has the desired dynamics. Indeed, we have
$\mathcal{K}^{\boldsymbol{\eta}}(0,t)^{\top}P_{t}=P_{0}-\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\big(P_{s}\widehat{A}_{s}+(\widehat{A}_{s})^{\top}P_{s}-P_{s}\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}P_{s}+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{s})\big)\mathrm{d}s$
$-\int_{0}^{t}(A^{1}_{s})^{\top}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}P_{s}\mathrm{d}\boldsymbol{\eta}_{s}$
$\mathcal{K}^{\boldsymbol{\eta}}(0,t)^{\top}\Pi_{t}=\Pi_{0}-\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\Big(\big(P_{s}+(\widehat{A}_{s})^{\top}-\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}\big)\Pi_{s}+P_{s}\widehat{D}^{(m,m^{\prime})}_{s}$
$+(\widehat{Q}_{s}+\widehat{\bar{Q}}_{s})\mathcal{K}^{\boldsymbol{\eta}}(0,s)M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}-\widehat{\bar{Q}}_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)S_{s}m_{s}\Big)\mathrm{d}s-\int_{0}^{t}(A^{1}_{s})^{\top}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\Pi_{s}\mathrm{d}\boldsymbol{\eta}_{s},$
and
$\mathcal{K}^{\boldsymbol{\eta}}(0,t)^{\top}P_{t}\mathcal{K}^{\boldsymbol{\eta}}(0,t)\big({X}^{(m,m^{\prime}),\xi}_{t}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{t}\big)=P_{0}\xi$
$-\int_{0}^{t}(A^{1}_{s})^{\top}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}P_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)\big({X}^{(m,m^{\prime}),\xi}_{s}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}$
$-\int_{0}^{t}\bigg[\Big(\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\big(P_{s}\widehat{A}_{s}+(\widehat{A}_{s})^{\top}P_{s}-P_{s}\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}P_{s}$
$+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{s})\big)\mathcal{K}^{\boldsymbol{\eta}}(0,s)\big({X}^{(m,m^{\prime}),\xi}_{s}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}\big)\Big)$
$-\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}P_{s}\Big(\widehat{A}_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)\big({X}^{(m,m^{\prime}),\xi}_{s}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}\big)$
$\quad\quad-\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}\mathcal{K}^{\boldsymbol{\eta}}(s,0)^{\top}Y^{(m,m^{\prime}),\xi}_{s}+\widehat{D}^{(m,m^{\prime})}_{s}\Big)\bigg]\,\mathrm{d}s$
$+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}P_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)\Sigma_{s}\mathrm{d}W_{s}.$
Keeping (38) in mind, we calculate the Riemann integral
$\int_{0}^{t}$ $\bigg[\Big(\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\big(P_{s}\widehat{A}_{s}+(\widehat{A}_{s})^{\top}P_{s}-P_{s}\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}P_{s}$
$+(\widehat{Q}_{t}+\widehat{\bar{Q}}_{s})\big)\mathcal{K}^{\boldsymbol{\eta}}(0,s)\big({X}^{(m,m^{\prime}),\xi}_{s}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}\big)\Big)$
$-\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}P_{s}\Big(\widehat{A}_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)\big({X}^{(m,m^{\prime}),\xi}_{s}-M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}\big)$
$-\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}\mathcal{K}^{\boldsymbol{\eta}}(s,0)^{\top}Y^{(m,m^{\prime}),\xi}_{s}+\widehat{D}^{(m,m^{\prime})}_{s}\Big)\bigg]\,\mathrm{d}s$
$+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\Big(\big(P_{s}+(\widehat{A}_{s})^{\top}-\widehat{B}_{s}(\widehat{R}_{s})^{-1}(\widehat{B}_{s})^{\top}\big)\Pi_{s}+P_{s}\widehat{D}^{(m,m^{\prime})}_{s}$
$+(\widehat{Q}_{s}+\widehat{\bar{Q}}_{s})\mathcal{K}^{\boldsymbol{\eta}}(0,s)M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{s}-\widehat{\bar{Q}}_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)S_{s}m_{s}\Big)\mathrm{d}s$
$=\int_{0}^{t}\big((Q_{s}+\bar{Q}_{s})X^{(m,m^{\prime}),\xi}_{s}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}\big)\mathrm{d}s.$
In view of the previous displays we deduce
$Y^{(m,m^{\prime}),\xi}_{t}=(Q+\bar{Q})X^{(m,m^{\prime}),\xi}_{T}-\bar{Q}Sm_{T}+\int_{t}^{T}(A^{1}_{s})^{\top}Y^{(m,m^{\prime}),\xi}_{s}\mathrm{d}\boldsymbol{\eta}_{s}-\int_{t}^{T}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}P_{s}\mathcal{K}^{\boldsymbol{\eta}}(0,s)\Sigma_{s}\mathrm{d}W_{s}$
$+\int_{t}^{T}\big((Q_{s}+\bar{Q}_{s})X^{(m,m^{\prime}),\xi}_{s}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{(m,m^{\prime}),\xi}_{s}\big)\mathrm{d}s.$
The existence proof is thus complete upon letting $Z^{(m,m^{\prime}),\xi}_{t}:=\mathcal{K}^{\boldsymbol{\eta}}(0,t)^{\top}P_{t}\mathcal{K}^{\boldsymbol{\eta}}(0,t)\Sigma_{t}.$ Additionally, we immediately get that
$(X^{(m,m^{\prime}),\xi})^{\prime}$ $=A^{1}{X}^{(m,m^{\prime}),\xi}+C^{1}m,$
$(Y^{(m,m^{\prime}),\xi})^{\prime}$ $=-(A^{1})^{\top}{Y}^{(m,m^{\prime}),\xi}.$
Turning to uniqueness, let
$\Big(\big(X^{(m,m^{\prime}),\xi},(X^{(m,m^{\prime}),\xi})^{\prime}\big),\big(Y^{(m,m^{\prime}),\xi},(Y^{(m,m^{\prime}),\xi})^{\prime}\big),Z^{(m,m^{\prime}),\xi}\Big)$
solve of (37) with the properties described in the statement of the theorem. By repeating the procedure, from the other direction, i.e. forward component first and then backward, with the help of A.3, we get that $\big(X^{(m,m^{\prime}),\xi},Y^{(m,m^{\prime}),\xi},\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot)^{\top}Z^{(m,m^{\prime}),\xi}\big)$ is a solution of (17). Notice that we are allowed to do that because, by definition, the solution space imposes sufficient integrability properties. In particular, $Z^{(m,m^{\prime}),\xi}$ is sufficiently integrable in order to apply A.3. Hence, from the uniqueness part of 3.9, uniqueness of (37) follows. ∎
4.3 Equivalence of rough fixed point step with FBRDE
To ease the notation and since the initial condition $\xi$ is fixed in this section, we shall omit it from the notation. Now, by taking expectations on (36) we get
$\left\{\begin{aligned} &\mathbb{E}\big[X^{(m,m^{\prime})}_{t}\big]=\mathcal{K}^{\boldsymbol{\eta}}(t,0)\mathbb{E}[\xi]+M^{\boldsymbol{\eta}}\big((m,m^{\prime})\big)_{t}\\ &\quad\quad\quad\quad\quad+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta}}(t,s)\big(A_{s}\mathbb{E}\big[X^{(m,m^{\prime})}_{s}\big]-B_{s}R_{s}^{-1}B_{s}^{\top}\mathbb{E}\big[Y^{(m,m^{\prime})}_{s}\big]+C_{s}m_{s}\big)\,\mathrm{d}s,\\ &\mathbb{E}\big[Y^{(m,m^{\prime})}_{t}\big]=\mathcal{K}^{\boldsymbol{\eta}}(T,t)^{\top}\big((Q+\bar{Q})\mathbb{E}\big[X^{(m,m^{\prime})}_{T}\big]-\bar{Q}Sm_{T}\big)\\ &\quad\quad\quad\quad\quad\quad+\int_{t}^{T}\mathcal{K}^{\boldsymbol{\eta}}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})\mathbb{E}\big[X^{(m,m^{\prime})}_{s}\big]-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathbb{E}\big[Y^{(m,m^{\prime})}_{s}\big]\Big)\mathrm{d}s.\end{aligned}\right$. (39)
From 4.5, with $\Sigma=0$ and $\xi$ deterministic (see the proof of 4.8), (39) is equivalent to the forward-backward RDE (FBRDE)
$\left\{\begin{aligned} &\mathbb{E}\big[X^{(m,m^{\prime})}_{t}\big]=\mathbb{E}[\xi]+\int_{0}^{t}\big(A^{1}_{s}\mathbb{E}\big[X^{(m,m^{\prime})}_{s}\big]+{C}^{(1)}_{s}m_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad\quad\quad\quad\quad+\int_{0}^{t}\big(A_{s}\mathbb{E}\big[X^{(m,m^{\prime})}_{s}\big]-B_{s}R_{s}^{-1}B_{s}^{\top}\mathbb{E}\big[Y^{(m,m^{\prime})}_{s}\big]+C_{s}m_{s}\big)\,\mathrm{d}s,\\ &\mathbb{E}\big[Y^{(m,m^{\prime})}_{t}\big]=(Q+\bar{Q})\mathbb{E}\big[X^{(m,m^{\prime})}_{T}\big]-\bar{Q}Sm_{T}+\int_{t}^{T}(A^{1}_{s})^{\top}\mathbb{E}\big[Y^{(m,m^{\prime})}_{s}\big]\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad\quad\quad\quad\quad+\int_{t}^{T}\Big((Q_{s}+\bar{Q}_{s})\mathbb{E}\big[X^{(m,m^{\prime})}_{s}\big]-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathbb{E}\big[Y^{(m,m^{\prime})}_{s}\big]\Big)\mathrm{d}s.\end{aligned}\right$. (40)
If $\big(m^{0},(m^{0})^{\prime}\big)$ is a fixed point of the map $\widetilde{F}:\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)\rightarrow\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$, where
$\widetilde{F}\big((m,m^{\prime})\big)_{t}:=\big(\mathbb{E}[{X}^{(m,m^{\prime})}_{t}],A^{1}_{t}\mathbb{E}[{X}^{(m,m^{\prime})}_{t}]+C^{1}_{t}m_{t}\big),$ (41)
then it satisfies the consistency equation
[…]
Plugging this equation to the system above and setting
$\mathbb{E}\big[X^{(m^{0},(m^{0})^{\prime})}_{t}\big]:=\bar{x}_{t},\mathbb{E}\big[Y^{(m^{0},(m^{0})^{\prime})}_{t}\big]:=\bar{y}_{t}$
we obtain
$\left\{\begin{aligned} &\bar{x}_{t}=\mathbb{E}[\xi]+\int_{0}^{t}(A^{1}_{s}+C^{1}_{s})\bar{x}_{s}\mathrm{d}\boldsymbol{\eta}_{s}+\int_{0}^{t}\big((A_{s}+C_{s})\bar{x}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}\bar{y}_{s}\big)\,\mathrm{d}s,\\ &\bar{y}_{t}=(Q+\bar{Q}-\bar{Q}S)\bar{x}_{T}+\int_{t}^{T}(A^{1}_{s})^{\top}\bar{y}_{s}\mathrm{d}\boldsymbol{\eta}_{s}+\int_{t}^{T}\big((Q_{s}+\bar{Q}_{s}-\bar{Q}_{s}S_{s})\bar{x}_{s}+A_{s}^{\top}\bar{y}_{s}\big)\mathrm{d}s.\end{aligned}\right$. (43)
Remark 4.9.
Note that the solution $\big((\bar{x},\bar{x}^{\prime}),(\bar{y},\bar{y}^{\prime})\big)$ of the FBRDE is an element of $\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)\times\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\bar{x}^{\prime}=(A^{1}+C^{1})\bar{x}$ and $\bar{y}^{\prime}=-(A^{1})^{\top}\bar{y}$.
Theorem 4.10.
There exists a one-to-one correspondence between the fixed points of map $\widetilde{F}$ (41) and the solutions of system (43). Specifically, every fixed point is the forward component of a solution of (43), and vise versa, every forward component of a solution of (43) is a fixed point of $\widetilde{F}$.
Proof.
Exactly the same as the proof of 4.2, with the usage of 4.5, or 4.8. ∎
4.4 Solving the FBRDE
From 4.10, the rough LQ MFG problem has a (unique) solution, provided that (43) has a (unique) solution.
In this section we solve the latter through a decoupling field $\bar{v}.$ After inspecting the form of the terminal condition in (43) we choose $\bar{v}$ to be linear. Thus we make the ansatz
$\bar{v}(t,x):=\bar{p}_{t}x,\hskip 2.84544pt\forall(t,x)\in[0,T]\times\mathbb{R}^{d}$
and
${\bar{y}}_{t}=\bar{p}_{t}{\bar{x}}_{t}.$ (44)
Of course, as we have already mentioned, in order for the decoupling field to be properly defined, we need to uniquely characterize the process $\bar{p}$. One way to do that is to find a proper characteristic equation for $\bar{p}$.
If $\boldsymbol{\eta}$ is smooth, we can absorb the rough dependence of (43) in the Riemann integrals, and the form of the characteristic equation for $\bar{p}$ is known by classical theory. This indicates that, in the rough case, the correct characteristic equation takes the form of the non-symmetric rough Riccati equation
$\bar{p}_{t}$ $={Q}+{\bar{Q}}-{\bar{Q}}{S}+\int_{t}^{T}\big(\bar{p}_{s}(A^{1}_{s}+{C}^{1}_{s})+(A^{1}_{s})^{\top}\bar{p}_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s}$ (45)
$+\int_{t}^{T}\Big(\bar{p}_{s}({A}_{s}+{C}_{s})+({A}_{s})^{\top}\bar{p}_{s}-\bar{p}_{s}{B}_{s}{R}_{s}^{-1}{B}_{s}^{\top}\bar{p}_{s}+{Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s}\Big)\mathrm{d}s.$
  1. (S4-R)
    We assume that $Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0$ and $Q+\bar{Q}-\bar{Q}S\succeq 0$. Additionally, let $(e,e^{\prime})\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d\times d}\big)$ be the unique solution to the linear RDE [^7] Uniqueness is standard, e.g. by straightforward extension of [31, Sec 8.9], or as special case of [12, Thm 3.11], or [42, Thm 3.5].
    $e_{t}=\text{Id}_{d\times d}+\int_{0}^{t}({C}_{s})^{\top}e_{s}\mathrm{d}s+\int_{0}^{t}({C}^{1}_{s})^{\top}e_{s}\mathrm{d}\boldsymbol{\eta}_{s}.$ (46)
    We assume that there exists $\lambda:[0,T]\rightarrow(0,\infty)$ such that, for all $t\in[0,T],$
    $e_{t}=\lambda_{t}\text{Id}_{d\times d}.$
    So, we get that $e^{\prime}_{t}=\lambda_{t}({C}^{1}_{t})^{\top}$.
Remark 4.11.
Assumption (S4-R) can be viewed as a natural ”rough” extension of an analogous condition that is typically assumed in classical LQ MFGs, see e.g. [19, Proposition 3.35]. It is satisfied when for example there exist $\lambda^{(1)}:[0,T]\rightarrow\mathbb{R}$, $\big(\lambda^{(2)},(\lambda^{(2)})^{\prime}\big)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{1\times p}\big)$, such that $C_{t}=\lambda^{(1)}_{t}\text{Id}_{d\times d}$ and ${C}^{1}_{t}(\cdot)=(\lambda^{(2)}_{t}\cdot)\text{Id}_{d\times d}$. Indeed then we have
$({C}_{t})^{\top}$ $=\lambda^{(1)}_{t}\text{Id}_{d\times d},$
$({C}^{1}_{t}(\cdot))^{\top}$ $=(\lambda^{(2)}_{t}\cdot)\text{Id}_{d\times d},$
and so $e_{t}=\exp{\Big(\int_{0}^{t}\lambda^{(1)}_{s}\mathrm{d}s+\int_{0}^{t}\lambda^{(2)}_{s}\mathrm{d}\boldsymbol{\eta}_{s}\Big)}\text{Id}_{d\times d}$.
Now, by A.2 we get a unique global solution for the symmetric rough Riccati
$\left\{\begin{aligned} &\widetilde{p}_{t}={(Q}+{\bar{Q}}-{\bar{Q}}{S})e^{-1}_{T}+\int_{t}^{T}\Big(\widetilde{p}_{s}(A^{1}_{s}+{C}^{1}_{s})+\big((A^{1}_{s})^{\top}+(C^{1}_{s})^{\top}\big)\widetilde{p}_{s}\Big)\mathrm{d}\boldsymbol{\eta}_{s}\\ &\quad\quad+\int_{t}^{T}\Big(\widetilde{p}_{s}({A}_{s}+{C}_{s})+({A}_{s})^{\top}\widetilde{p}_{s}-\widetilde{p}_{s}{B}_{s}({R}_{s}e^{-1}_{s})^{-1}{B}_{s}^{\top}\widetilde{p}_{s}+({Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s})e^{-1}_{s}\Big)\mathrm{d}s.\end{aligned}\right$. (47)
We define, $\bar{p}_{t}:=e_{t}\widetilde{p}_{t}$ and we have
$({C}^{1}_{t})^{\top}e_{t}\widetilde{p}_{t}-e_{t}\big(\widetilde{p}_{t}(A^{1}_{t}+{C}^{1}_{t})+\big(({A}^{1}_{t})^{\top}+({C}^{1}_{t})^{\top}\big)\widetilde{p}_{t}\big)=-\big(\bar{p}_{t}(A^{1}_{t}+{C}^{1}_{t})+(A^{1}_{t})^{\top}\bar{p}_{t}\big),$
$({C}_{t})^{\top}e_{t}\widetilde{p}_{t}-e_{t}\big(\widetilde{p}_{t}({A}_{t}+{C}_{t})+({A}_{t}+{C}_{t})^{\top}\widetilde{p}_{t}-\widetilde{p}_{t}{B}_{t}({R}_{t}e_{t}^{-1})^{-1}{B}_{t}^{\top}\widetilde{p}_{t}$
$+({Q}_{t}+{\bar{Q}}_{t}-{\bar{Q}}_{t}{S}_{t})e_{t}^{-1}\big)$
$=-\big(\bar{p}_{t}({A}_{t}+{C}_{t})+({A}_{t})^{\top}\bar{p}_{t}-\bar{p}_{t}{B}_{t}{R}_{t}^{-1}{B}_{t}^{\top}\bar{p}_{t}+{Q}_{t}+{\bar{Q}}_{t}-{\bar{Q}}_{t}{S}_{t}\big).$
Then, by the rough product formula, A.3, we get that $\bar{p}_{t}$ satisfies (45). So, $\bar{v}(t,x)$ is well defined [^8] Uniqueness of (45) comes by iterating local uniqueness.. Then, using the ansatz we decouple the first equation of (43) and write it as
$\bar{x}_{t}=\mathbb{E}[\xi]+\int_{0}^{t}(A^{1}_{s}+C^{1}_{s})\bar{x}_{s}\mathrm{d}\boldsymbol{\eta}_{s}+\int_{0}^{t}(A_{s}+C_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}\bar{p}_{s})\bar{x}_{s}\,\mathrm{d}s.$ (48)
As a linear RDE, equation (48) has a unique global solution in $\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$. Finally, we define
${\bar{y}}_{t}:=\bar{v}(t,\bar{x}_{t})=\bar{p}_{t}{\bar{x}}_{t},$
where ${\bar{x}}_{t}$ solves (48). To finish with existence, we need to verify that it satisfies the backward equation
${\bar{y}}_{t}=({Q}+{\bar{Q}}-{\bar{Q}}{S}){\bar{x}}_{T}+\int_{t}^{T}(A^{1}_{s})^{\top}\bar{p}_{s}\bar{x}_{s}\mathrm{d}\boldsymbol{\eta}_{s}+\int_{t}^{T}\Big({Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s}+({A}_{s})^{\top}\bar{p}_{s}\Big){\bar{x}}_{s}\mathrm{d}s,$
in other words, we want to holds
$\bar{p}_{t}{\bar{x}}_{t}$ $=({Q}+{\bar{Q}}-{\bar{Q}}{S}){\bar{x}}_{T}+\int_{t}^{T}(A^{1}_{s})^{\top}\bar{p}_{s}\bar{x}_{s}\mathrm{d}\boldsymbol{\eta}_{s}+\int_{t}^{T}\Big({Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s}+{A}_{s}^{\top}\bar{p}_{s}\Big){\bar{x}}_{s}\mathrm{d}s$
$=\bar{p}_{0}\mathbb{E}[\xi]-\int_{0}^{t}(A^{1}_{s})^{\top}\bar{p}_{s}\bar{x}_{s}\mathrm{d}\boldsymbol{\eta}_{s}-\int_{0}^{t}\Big({Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s}+{A}_{s}^{\top}\bar{p}_{s}\Big){\bar{x}}_{s}\mathrm{d}s,$
where $\bar{p}_{t}$ satisfies equation (45) and $\bar{x}_{t}$ satisfies (48). This can be proven immediately by using the rough product formula A.3 as above. Indeed, we have
$-\Big(\bar{p}_{t}({A}_{t}+{C}_{t})+{A}_{t}^{\top}\bar{p}_{t}-\bar{p}_{t}{B}_{t}{R}_{t}^{-1}{B}_{t}^{\top}\bar{p}_{t}+({Q}_{t}+{\bar{Q}}_{t}-{\bar{Q}}_{t}{S}_{t})\Big){\bar{x}}_{t}$
$+\bar{p}_{t}\big({A}_{t}+{C}_{t}-{B}_{t}{R}_{t}^{-1}{B}_{t}^{\top}\bar{p}_{t}\big){\bar{x}}_{t}$
$=-\Big({A}_{t}^{\top}\bar{p}_{t}+({Q}_{t}+{\bar{Q}}_{t}-{\bar{Q}}_{t}{S}_{t})\Big){\bar{x}}_{t}$
and so the proof for existence of a solution of (43) is complete.
Next, we prove uniqueness. To this end we define $\Xi^{\boldsymbol{\eta}}_{t\leftarrow s},\Xi^{\boldsymbol{\eta}}_{s\leftarrow t}$ as the unique solutions of
$\Xi^{\boldsymbol{\eta}}_{t\leftarrow s}=\text{Id}_{d\times d}+\int_{s}^{t}(A^{1}_{r}+{C}^{1}_{r})\Xi^{\boldsymbol{\eta}}_{r\leftarrow s}\mathrm{d}\boldsymbol{\eta}_{r},$
$\Xi^{\boldsymbol{\eta}}_{s\leftarrow t}=\text{Id}_{d\times d}-\int_{s}^{t}\Xi^{\boldsymbol{\eta}}_{s\leftarrow r}(A^{1}_{r}+{C}^{1}_{r})\mathrm{d}\boldsymbol{\eta}_{s}.$
Additionally, we define
$\Lambda^{\boldsymbol{\eta}}(t,s):=\Xi^{\boldsymbol{\eta}}_{t\leftarrow 0}\Xi^{\boldsymbol{\eta}}_{0\leftarrow s}.$
By virtue of A.3, and by repeating the proof of 4.8 (without worrying about integrability properties as we are in a deterministic setting), we see that (43) is equivalent to
$\left\{\begin{aligned} {\bar{x}}_{t}&=\Lambda^{\boldsymbol{\eta}}(t,0)\mathbb{E}[\xi]+\int_{0}^{t}\Lambda^{\boldsymbol{\eta}}(t,s)\big(({A}_{s}+{C}_{s}){\bar{x}}_{s}-{B}_{s}{R}_{s}^{-1}{B}_{s}^{\top}{\bar{y}}_{s}\big)\,\mathrm{d}s,\\ {\bar{y}}_{t}&=\mathcal{K}^{\boldsymbol{\eta}}(T,t)^{\top}({Q}+{\bar{Q}}-{\bar{Q}}{S}){\bar{x}}_{T}+\int_{t}^{T}\mathcal{K}^{\boldsymbol{\eta}}(s,t)^{\top}\Big(({Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s}){\bar{x}}_{s}+({A}_{s})^{\top}{\bar{y}}_{s}\Big)\mathrm{d}s,\end{aligned}\right$. (49)
and by defining $\widehat{\bar{x}}_{t}:=\Lambda^{\boldsymbol{\eta}}(0,t)\bar{x}_{t},\hskip 2.84544pt\widehat{\bar{y}}_{t}:=\mathcal{K}^{\boldsymbol{\eta}}(t,0)^{\top}\bar{y}_{t}$, system (49) is equivalent to
$\left\{\begin{aligned} &\widehat{\bar{x}}_{t}=\mathbb{E}[\xi]+\int_{0}^{t}\Lambda^{\boldsymbol{\eta}}(0,s)\big(({A}_{s}+{C}_{s})\Lambda^{\boldsymbol{\eta}}(s,0)\widehat{\bar{x}}_{s}-{B}_{s}{R}_{s}^{-1}{B}_{s}^{\top}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\widehat{\bar{y}}_{s}\big)\,\mathrm{d}s,\\ &\widehat{\bar{y}}_{t}=\mathcal{K}^{\boldsymbol{\eta}}(T,0)^{\top}({Q}+{\bar{Q}}-{\bar{Q}}{S})\Lambda^{\boldsymbol{\eta}}(T,0)\widehat{\bar{x}}_{T}\\ &\quad\quad+\int_{t}^{T}\mathcal{K}^{\boldsymbol{\eta}}(s,0)^{\top}\Big(({Q}_{s}+{\bar{Q}}_{s}-{\bar{Q}}_{s}{S}_{s})\Lambda^{\boldsymbol{\eta}}(s,0)\widehat{\bar{x}}_{s}+({A}_{s})^{\top}\mathcal{K}^{\boldsymbol{\eta}}(0,s)^{\top}\widehat{\bar{y}}_{s}\Big)\mathrm{d}s.\end{aligned}\right$. (50)
From the analysis at the beginning of the subsection we have shown that (43) has a solution of the form $(\bar{x}_{t},\bar{p}_{t}\bar{x}_{t})$. Hence, system (50) has a solution in the form $\big(\widehat{\bar{x}}_{t},\mathcal{K}^{\boldsymbol{\eta}}(t,0)^{\top}\bar{p}_{t}\Lambda^{\boldsymbol{\eta}}(t,0)\widehat{\bar{x}}_{t}\big)$. In other words, there exists a decoupling field that gives a solution of (50), so, from [19, Proposition 4.8] this solution is unique. Thus, going backwards, (43) also has a unique solution, and we have arrived at the following result.
Theorem 4.12.
Let (S1-R), (S2), (S3) and (S4-R) hold true. Then, (43) has a unique solution, and equivalently, the rough LQ MFG also has a unique solution.
Remark 4.13.
(50) is a standard FBODE that is equivalent to (43) and one may attempt to solve it directly. However, in order to deal with the asymmetric terms, existence of global solutions would require an additional condition of the form:
  1. (S5-R)
    $Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0$ and $Q+\bar{Q}-\bar{Q}S\succeq 0$. Additionally, with $e:[0,T]\rightarrow\mathbb{R}^{d\times d}$ unique solution of
    $e_{t}=\text{Id}_{d\times d}+\int_{0}^{t}(\widehat{C}_{s})^{\top}e_{s}\mathrm{d}s,$
    and $\widehat{C}_{t}:=\mathcal{K}^{\boldsymbol{\eta}}(0,t)C_{t}\mathcal{K}^{\boldsymbol{\eta}}(t,0),$ there exist $\lambda^{(1)},\lambda^{(2)}:[0,T]\rightarrow(0,\infty)$ such that for all $t\in[0,T],$
    $\Lambda^{\boldsymbol{\eta}}(t,0)$ $=\lambda^{(1)}_{t}\mathcal{K}^{\boldsymbol{\eta}}(t,0),$
    $e_{t}$ $=\lambda^{(2)}_{t}\text{Id}_{d\times d}.$
5 Stability of the rough LQ MFG
Let $p\in\mathbb{N},\gamma\in(\frac{1}{3},\frac{1}{2})$ and $\mathcal{A}$ as in (30). In this section we establish stability of the solution to the MFG problem (1)-(3) with respect to initial conditions and the rough common noise. In particular, we prove (local Lipschitz) continuity properties for the equilibrium law, optimal value function and optimal control, respectively given by
$\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})\times\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d})\ni(\boldsymbol{\eta},\xi)\longmapsto\mu^{\boldsymbol{\eta},\xi}=\mathcal{L}(X^{\boldsymbol{\eta},\alpha^{*}(\boldsymbol{\eta},\xi),\xi}_{\cdot})\in\mathcal{W}_{k}(C^{\gamma}([0,T];\mathbb{R}^{d})),\;k\in[2,\infty),$
$\alpha^{*}(\boldsymbol{\eta},\xi)=\mathrm{argmin}_{\alpha\in\mathcal{A}}\,\mathcal{J}(\alpha\,;\boldsymbol{\eta},(m,m^{\prime})(\boldsymbol{\eta},\mathbb{E}[\xi]),\xi),$
and
$\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})\ni\boldsymbol{\eta}\longmapsto\mathcal{J}(\alpha^{*};\boldsymbol{\eta},(m,m^{\prime}),\xi)\in\mathbb{R}.$
Above and throughout this section $\mathcal{W}_{k}(\mathcal{X}),k\in[2,\infty),$ denotes the $k$-Wasserstein space of probability measures over a Polish space $\mathcal{X}$ with topology induced by a metric $d.$ For $\mu^{1},\mu^{2}\in\mathcal{W}_{k}(\mathcal{X})$ the $k$-Wasserstein distance is given by $\mathcal{W}^{k}_{k,\mathcal{X}}(\mu^{1},\mu^{2})=\inf_{\pi\in\Pi(\mu^{1},\mu^{2})}\iint_{\mathcal{X}\times\mathcal{X}}d^{k}(x,y)\mathrm{d}\pi(x,y),$ where $\Pi$ denotes the family of couplings between $\mu^{1},\mu^{2}.$
Remark 5.1.
As we have already mentioned, the kernel $\mathcal{K}^{\boldsymbol{\eta}}$ and initial path $M^{\boldsymbol{\eta}}$ depend on $(A^{1},(A^{1})^{\prime})$ and $(A^{1},(A^{1})^{\prime}),(C^{1},(C^{1})^{\prime})$ respectively. For ease on notation, we continue to omit these dependencies. Nevertheless, they are implicitly assumed and are, in fact, important for the following stabilility analysis.
The following auxiliary lemma establishes continuity properties of the mean flows $(m,m^{\prime})$ with respect to initial conditions and common noise.
Lemma 5.2.
Let $\xi_{1},\xi_{2}\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d}),T>0,$ $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. For $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$ that satisfy Assumption (S4-R) let $(m_{i},(A^{1,i}+C^{1,i})m_{i})\in\mathscr{D}^{\gamma,\beta}_{\eta_{i}}\big([0,T];\mathbb{R}^{d}\big)$ solve (43) with coefficients $(A^{1,i},(A^{1,i})^{\prime}),(C^{1,i},(C^{1,i})^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta_{i}}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ and initial conditions $\mathbb{E}[\xi_{i}],$ $i=1,2.$ Under Assumptions (S1-R), (S2), (S3), if $M>0$ is a constant such that
$\rho_{\gamma}(\boldsymbol{\eta}_{1})+\rho_{\gamma}(\boldsymbol{\eta}_{2})\leq M$
then, up to a nonnegative multiplicative constant that depends on $M,T,\gamma,\beta,$
$\|m_{1}-m_{2}\|_{\infty;[0,T]}$ $+\|(m_{1},(A^{1,1}+C^{1,1})m_{1});(m_{2},(A^{1,2}+C^{1,2})m_{2})\|_{\eta_{1},\eta_{2},\gamma,\beta}$
$\lesssim\|\xi_{2}-\xi_{1}\|_{k}+\rho_{\gamma}(\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}.$
Proof.
In view of the consistency condition (42), any solution $(m,m^{\prime})$ of (43) satisfies
$({m},{m}^{\prime})=(\mathbb{E}[X^{\alpha^{*}}],\mathbb{E}[(X^{\alpha^{*}})^{\prime}])=(\bar{x},\bar{x}^{\prime})=(\bar{x},(A^{1}+C^{1})\bar{x})$
on $[0,T]$ and $(\bar{x},\bar{x}^{\prime})\in\mathscr{D}^{\gamma,\beta}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ is the forward component of the FBRDE (43). From Theorem 4.12, the unique solution $(\bar{x},\bar{y})$ of the latter satisfies the ansatz (44). Consequently, we can write the solution pair in the form $(\bar{x},\bar{p}\bar{x}),$ where $\bar{p}$ solves the (non-symmetric) rough Riccati equation (45). In turn, as shown in Section 4.4, $\bar{p}=e\tilde{p}$, where $e$ is the (unique, global) solution of the linear RDE (46) and $\tilde{p}$ solves (uniquely and globally) the symmetric rough Riccati (47). Thus, $(\bar{x},\bar{y})$ takes the form $(\bar{x},e\tilde{p}\bar{x})$ and, in view of Theorem A.2, the maps $\boldsymbol{\eta}\longmapsto e,\tilde{p}$ are locally Lipschitz continuous. Returning to the FBRDE, we substitute $\bar{y}=e\hat{p}\bar{x}$ to the dynamics of the forward component and obtain the decoupled linear RDE
$\bar{x}_{t}=\mathbb{E}[\xi]+\int_{0}^{t}(A^{1}_{s}+C^{1}_{s})\bar{x}_{s}\mathrm{d}\boldsymbol{\eta}_{s}+\int_{0}^{t}\big((A_{s}+C_{s})\bar{x}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}e_{s}\tilde{p}_{s}\bar{x}_{s}\big)\,\mathrm{d}s.$ (51)
To obtain the desired stability estimate, let $(\bar{x}_{i},\bar{x}^{\prime}_{i})$ and $\bar{p}^{i}=e^{i}\tilde{p}^{i},i=1,2,$ denote respectively solutions to (51) and (45) driven by $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$ and with initial conditions $\xi_{1},\xi_{2}.$ Applying a stability estimate for (linear, deterministic) RDEs (see Proposition A.1), and noting that the drift terms also depend on the rough drivers, we obtain
$\|\bar{x}_{1}-\bar{x}_{2}\|_{\infty;[0,T]}$ $+\|(\bar{x}_{1},\bar{x}^{\prime}_{1});(\bar{x}_{2},\bar{x}^{\prime}_{2})\|_{\eta_{1},\eta_{2},\gamma,\beta}$
$\lesssim|\mathbb{E}[\xi_{2}]-\mathbb{E}[\xi_{1}]|+\rho_{\gamma}(\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2})$
$+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\big\|e^{2}\tilde{p}^{2}-e^{1}\tilde{p}^{1}\big\|_{\infty;[0,T]},$
[…]
$\big\|e^{2}\tilde{p}^{2}-e^{1}\tilde{p}^{1}\big\|_{\infty;[0,T]}$ $\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}.$
The proof is complete. ∎
We turn to the proof of our main stability result, Theorem 1.4 which we state in detail below.
Theorem 5.3 (Stability of the rough MFG solution).
Let $\xi_{1},\xi_{2}\in\bigcap_{k=2}^{\infty}\mathbb{L}^{k}(\mathcal{F}_{0},\mathbb{R}^{d}),T>0,$ $\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. For $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$ that satisfy Assumption (S4-R) let $(m_{i},m^{\prime}_{i})\in\mathscr{D}^{\gamma,\beta}_{\eta_{i}}\big([0,T];\mathbb{R}^{d}\big)$ solve (43) with coefficients $(A^{1,i},(A^{1,i})^{\prime}),$ $(C^{1,i},(C^{1,i})^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta_{i}}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ and initial conditions $\mathbb{E}[\xi_{i}]$ and $(X^{\boldsymbol{\eta}_{i},\alpha^{*,i},(m_{i},m_{i}^{\prime})},\alpha^{*,i}),$ $i=1,2$ be the corresponding solutions to the MFG (1)-(3). Under Assumptions (S1-R), (S2), (S3), if $M>0$ is a constant such that
$\rho_{\gamma}(\boldsymbol{\eta}_{1})+\rho_{\gamma}(\boldsymbol{\eta}_{2})\leq M$
then, up to a nonnegative multiplicative constant that depends on $M,T,\gamma,\beta,\xi,$ we have
$\sup_{t\in[0,T]}$ $\|\alpha^{*,2}_{t}-\alpha^{*,1}_{t}\|_{k}+\|\|X^{2}-X^{1}\|_{\gamma}\|_{k}+\mathcal{W}_{k,C^{\gamma}}(\mu^{1},\mu^{2})$ (52)
$+\big|\mathcal{J}(\alpha^{*,2};\boldsymbol{\eta}_{2},(m_{2},m^{\prime}_{2}),\xi_{2})-\mathcal{J}(\alpha^{*,1};\boldsymbol{\eta}_{1},(m_{1},m_{1}^{\prime}),\xi_{1})\big|$
$\lesssim\|\xi_{2}-\xi_{1}\|_{k}+\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}},$
where $\mu^{i}=\mathcal{L}(X^{i})=\mathcal{L}(X^{\boldsymbol{\eta}_{i},\alpha^{*,i},(m_{i},m^{\prime}_{i}),\xi_{i}})\in\mathcal{W}_{k}(C^{\gamma}([0,T];\mathbb{R}^{d}),$ and the optimal controls
$\alpha^{*,i}=\mathrm{argmin}_{\alpha\in\mathcal{A}}\,\mathcal{J}(\alpha\,;\boldsymbol{\eta}_{i},(m_{i},m_{i}^{\prime}),\xi_{i}),\;\;i=1,2$
take the explicit form $\alpha^{*,i}_{t}=-R_{t}^{-1}B_{t}^{\top}Y^{i}_{t},t\in[0,T]$ with $Y^{i}$ solving (37). Finally, if the controlled rough paths $A^{1,i},C^{1,i},i=1,2$ are such that
$\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1}),$ (53)
we obtain the continuity estimate
$\sup_{t\in[0,T]}$ $\|\alpha^{*,2}_{t}-\alpha^{*,1}_{t}\|_{k}+\|\|X^{2}-X^{1}\|_{\gamma}\|_{k}+\mathcal{W}_{k,C^{\gamma}}(\mu^{1},\mu^{2})$
$+\big|\mathcal{J}(\alpha^{*,2};\boldsymbol{\eta}_{2},(m_{2},m_{2}^{\prime}),\xi_{2})-\mathcal{J}(\alpha^{*,1};\boldsymbol{\eta}_{1},(m_{1},m_{1}^{\prime}),\xi_{1})\big|\lesssim\|\xi_{2}-\xi_{1}\|_{k}+\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1}).$
Proof.
Starting from the optimal state, we adopt the shorter notation
$X^{i}=X^{\boldsymbol{\eta}_{i},\alpha^{*,i},(m_{i},m^{\prime}_{i}),\xi_{i}},i=1,2,$
with $\alpha^{*,i},(m_{i},m^{\prime}_{i})$ denoting the optimal control and mean paths of the MFG with corresponding common rough noise $\boldsymbol{\eta}_{i}.$ In view of Theorem 1.1, we express the dynamics in Volterra form and, for $0\leq s<t\leq T,$
$\delta X^{i}_{s,t}\equiv X^{i}_{t}$ $-X^{i}_{s}=\bigg(\mathcal{K}^{\boldsymbol{\eta_{i}}}(t,0)-\mathcal{K}^{\boldsymbol{\eta_{i}}}(s,0)\bigg)\xi_{i}+M^{\boldsymbol{\eta}_{i}}\big((m_{i},m^{\prime}_{i})\big)_{t}-M^{\boldsymbol{\eta}_{i}}\big((m_{i},m^{\prime}_{i})\big)_{s}$ (54)
$+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta_{i}}}(t,r)\bigg(A_{r}X^{i}_{r}+\alpha^{*,i}_{r}+C_{r}m^{i}_{r}\bigg)\mathrm{d}r-\int_{0}^{s}\mathcal{K}^{\boldsymbol{\eta_{i}}}(s,r)\bigg(A_{r}X^{i}_{r}+\alpha^{*,1}_{r}+C_{r}m^{i}_{r}\bigg)\mathrm{d}r$
$+\int_{0}^{t}\mathcal{K}^{\boldsymbol{\eta_{i}}}(t,r)\Sigma_{r}\mathrm{d}W_{r}-\int_{0}^{s}\mathcal{K}^{\boldsymbol{\eta_{i}}}(s,r)\Sigma_{r}\mathrm{d}W_{r}$
$=:\delta(\mathcal{K}^{\boldsymbol{\eta_{i}}})_{s,t}\xi_{i}+\delta M^{\boldsymbol{\eta}_{i}}\big((m_{i},m^{\prime}_{i})\big)_{s,t}$
$+\int_{s}^{t}\mathcal{K}^{\boldsymbol{\eta_{i}}}(t,r)\bigg(A_{r}X^{i}_{r}+\alpha^{*,i}_{r}+C_{r}m^{i}_{r}\bigg)\mathrm{d}r$
$+\delta(\mathcal{K}^{\boldsymbol{\eta_{i}}})_{s,t}\int_{0}^{s}\mathcal{K}^{\boldsymbol{\eta_{i}}}(0,r)\bigg(A_{r}X^{i}_{r}+\alpha^{*,i}_{r}+C_{r}m^{i}_{r}\bigg)\mathrm{d}r$
$+\int_{s}^{t}\mathcal{K}^{\boldsymbol{\eta_{i}}}(t,r)\Sigma_{r}\mathrm{d}W_{r}+\delta(\mathcal{K}^{\boldsymbol{\eta_{i}}})_{s,t}\int_{0}^{s}\mathcal{K}^{\boldsymbol{\eta_{i}}}(0,r)\Sigma_{r}\mathrm{d}W_{r}.$
Note that the first two terms are deterministic and state-independent. Since the Volterra kernel (33) solves the linear matrix-valued RDE (31) and $M^{\eta_{i}}$ is given by the (deterministic) rough integral (35) we have the estimates
$\|\delta(\mathcal{K}^{\boldsymbol{\eta_{2}}})\xi_{2}-\delta(\mathcal{K}^{\boldsymbol{\eta_{1}}})\xi_{1}\|_{\gamma}\lesssim|\xi_{2}-\xi_{1}|+\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}},$ (55)
$\|\delta M^{\boldsymbol{\eta}_{2}}\big((m_{2},m^{\prime}_{2})\big)-\delta M^{\boldsymbol{\eta}_{1}}\big((m_{1},m^{\prime}_{1})\big)\|_{\gamma}$ (56)
$\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\bigg\|\int_{0}^{\cdot}\big(Z_{2}^{\boldsymbol{\eta_{2}},(m_{2},m_{2}^{\prime})},(Z_{2}^{\prime})^{\boldsymbol{\eta}_{2},(m_{2},m_{2}^{\prime})}\big)_{s}\mathrm{d}\boldsymbol{\eta}_{2,s}-\int_{0}^{\cdot}\big(Z_{1}^{\boldsymbol{\eta_{1}},(m_{1},m_{1}^{\prime})},(Z_{1}^{\prime})^{\boldsymbol{\eta}_{1},(m_{1},m_{1}^{\prime})}\big)_{s}\mathrm{d}\boldsymbol{\eta}_{1,s}\bigg\|_{\gamma}.$
By classical stability estimates for rough integrals (see e.g. [31, Chapter 4]), the last term is bounded from above by
$\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\big\|\big(Z_{2}^{\boldsymbol{\eta_{2}},(m_{2},m_{2}^{\prime})},(Z_{2}^{\prime})^{\boldsymbol{\eta}_{2},(m_{2},m_{2}^{\prime})}\big);\big(Z_{1}^{\boldsymbol{\eta_{1}},(m_{1},m_{1}^{\prime})},(Z_{1}^{\prime})^{\boldsymbol{\eta}_{1},(m_{1},m_{1}^{\prime})}\big)\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}.$
In turn, in view of Definition (4.6), the second term above is bounded from above by
$\big\|$ $(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}+\big\|\big(m_{2},m_{2}^{\prime}\big);\big(m_{1},m_{1}^{\prime}\big)\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\big\|\big(\mathcal{K}^{\boldsymbol{\eta_{2}}}(0,\cdot),(\mathcal{K}^{\boldsymbol{\eta_{2}}})^{\prime}(0,\cdot)\big);\big(\mathcal{K}^{\boldsymbol{\eta_{1}}}(0,\cdot),(\mathcal{K}^{\boldsymbol{\eta_{1}}})^{\prime}(0,\cdot)\big)\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}.$
We complete the estimate for the second term by invoking Lemma 5.2 along with stability estimates for the linear RDE (31). These bounds then furnish
$\|\delta M^{\boldsymbol{\eta}_{2}}\big((m_{2},m^{\prime}_{2})\big)$ $-\delta M^{\boldsymbol{\eta}_{1}}\big((m_{1},m^{\prime}_{1})\big)\|_{\gamma}$ (57)
$\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\big\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\big\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}.$
Turning to the third term in (54) we have
$\bigg|\int_{s}^{t}\mathcal{K}^{\boldsymbol{\eta_{2}}}$ $(t,r)\bigg(A_{r}X^{2}_{r}+\alpha^{*,2}_{r}+C_{r}m^{2}_{r}\bigg)\mathrm{d}r-\int_{s}^{t}\mathcal{K}^{\boldsymbol{\eta_{1}}}(t,r)\bigg(A_{r}X^{1}_{r}+\alpha^{*,1}_{r}+C_{r}m^{1}_{r}\bigg)\mathrm{d}r\bigg|$ (58)
$\lesssim\|\delta(\mathcal{K}^{\boldsymbol{\eta_{2}}})-\delta(\mathcal{K}^{\boldsymbol{\eta_{1}}})\|_{\gamma}\int_{s}^{t}(t-r)^{\gamma}\bigg(1+|\alpha^{*,2}_{r}|+|m_{r}^{2}|\bigg)\mathrm{d}r$
$+\int_{s}^{t}\bigg(|A_{r}||X^{1}_{r}-X^{2}_{r}|+|\alpha^{*,2}_{r}-\alpha^{*,1}_{r}|+|m^{2}_{r}-m^{1}_{r}|\bigg)\mathrm{d}r$
$\lesssim\bigg[\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime};[0,T]}\bigg](t-s)^{\gamma}\int_{0}^{T}\bigg(1+|\alpha^{*,2}_{r}|\bigg)\mathrm{d}r$
$+\int_{0}^{T}|X^{1}_{r}-X^{2}_{r}|\mathrm{d}r+\int_{s}^{t}|\alpha^{*,2}_{r}-\alpha^{*,1}_{r}|\mathrm{d}r+(t-s)\|m^{2}-m^{1}\|_{\infty;[0,T]},$
almost surely, where we used (55) as well as standard estimates for $\|m^{2}\|_{\infty;[0,T]},\|U^{\boldsymbol{\eta}_{1}}_{0\leftarrow\cdot}\|_{\infty;[0,T]}$ which are absorbed in the multiplicative constant, to obtain the fourth line. Furthermore notice that, by virtue of Lemma (5.2) the last term is upper bounded by
$(t-s)\bigg(\|\xi_{2}-\xi_{1}\|_{k}+$ $\rho_{\gamma}(\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\bigg).$
In order to proceed, we need an estimate on the optimal controls. To this end, recall from (36) that
$\alpha^{*,i}_{t}=-R_{t}^{-1}B_{t}^{\top}Y^{i}_{t}$
and the unique, global solution $Y^{i}$ of the backward RDE satisfies the ansatz (38)
$Y_{t}^{i}=\mathcal{K}^{\boldsymbol{\eta}_{i}}(0,t)^{\top}P^{\boldsymbol{\eta}_{i}}_{t}\mathcal{K}^{\boldsymbol{\eta}_{i}}(0,t)\big({X}^{i}_{t}-M^{\boldsymbol{\eta}_{i}}\big((m_{i},m_{i}^{\prime})\big)_{t}\big)+\mathcal{K}^{\boldsymbol{\eta}_{i}}(0,t)^{\top}\Pi^{\boldsymbol{\eta}_{i},(m_{i},m_{i}^{\prime})}_{t}$.
Thus, since $R,B$ are deterministic and bounded uniformly over $t\in[0,T],$ we have
$|\alpha^{*,2}_{r}-\alpha^{*,1}_{r}|\lesssim|Y^{2}_{r}-Y^{1}_{r}|$
almost surely for all $r\in[0,t].$ Next, notice that $P^{\boldsymbol{\eta}_{i}},\Pi^{\boldsymbol{\eta}_{i},(m_{i},m_{i}^{\prime})}$ are (differentiable-in-time) solutions of the deterministic system of backward matrix equations (27) with coefficients $\widehat{A}^{i},\widehat{B}^{i},\widehat{Q}^{i},\widehat{\bar{Q}}^{i},\mathcal{K}^{\boldsymbol{\eta}_{i}}$ depending on $\boldsymbol{\eta}_{i},i=1,2.$ Thus, up to multiplicative constants that depend polynomially on the seminorms $\rho_{\gamma}(\boldsymbol{\eta}_{i})$ we have the almost-sure estimate
$|\alpha^{*,2}_{r}-\alpha^{*,1}_{r}|$ $\lesssim\|\xi_{2}-\xi_{1}\|_{k}+\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\big\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$ (59)
$+\big\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}+|X^{2}_{r}-X^{1}_{r}|.$
Substituting the latter back to (58) we obtain, up to deterministic constants,
$\bigg|\int_{s}^{t}\mathcal{K}^{\boldsymbol{\eta_{2}}}(t,r)\bigg(A_{r}X^{2}_{r}+\alpha^{*,2}_{r}+C_{r}m^{2}_{r}\bigg)\mathrm{d}r-\int_{s}^{t}\mathcal{K}^{\boldsymbol{\eta_{1}}}(t,r)\bigg(A_{r}X^{1}_{r}+\alpha^{*,1}_{r}+C_{r}m^{1}_{r}\bigg)\mathrm{d}r\bigg|$ (60)
$\lesssim(t-s)^{\gamma}\bigg(\|\xi_{2}-\xi_{1}\|_{k}+\rho_{\gamma}(\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$
$+\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\bigg)+\int_{0}^{T}|X^{2}_{r}-X^{1}_{r}|\mathrm{d}r$
and an identical pathwise bound holds for the fourth term in (54); its proof is omitted.
It remains to estimate the Itô integral terms in (54). Both can be treated similarly via Doob’s inequality. For example, for the first term on the last line we obtain, for any $k\geq 1$
$\bigg\|\int_{s}^{t}(\mathcal{K}^{\boldsymbol{\eta_{2}}}$ $(t,r)-\mathcal{K}^{\boldsymbol{\eta_{2}}}(t,r))\Sigma_{r}\mathrm{d}W_{r}\bigg\|_{k}$ (61)
$\lesssim(t-s)^{1/2}\bigg(\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\bigg).$
Collecting the estimates (56), (57), (60), (61) and applying a standard Grönwall inequality yields the following stability bound for the optimal state:
$\|\|X^{2}-X^{1}\|_{\gamma}\|_{k}$ $\lesssim\|\xi_{2}-\xi_{1}\|_{k}+\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1})+\big\|(C^{1,1},(C^{1,1})^{\prime});(C^{1,2},(C^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}$ (62)
$+\big\|(A^{1,1},(A^{1,1})^{\prime});(A^{1,2},(A^{1,2})^{\prime})\big\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}.$
Returning to (59) we obtain the same bound for
$\sup_{t\in[0,T]}\|\alpha^{*,2}_{t}-\alpha^{*,1}_{t}\|_{k}.$
The estimate for the optimal costs $J(\alpha^{i,*},\xi_{i})$ follows immediately from the previous estimates and the fact that the cost function is quadratic in the state, mean vectors and controls. Finally, we consider the $k$-Wasserstein distance between the equilibrium laws $\mu^{i},i=1,2,$ viewed as measures on the Hölder space $C^{\gamma}([0,T];\mathbb{R}^{d})$ (with norm $\|\cdot\|_{\gamma}$ defined by the sum of supremum norm and Hölder seminorm). The desired bound follows from (62) along with the straightforward inequality
$\mathcal{W}_{k,C^{\gamma}}(\mu^{1},\mu^{2})^{k}\leq\mathbb{E}[\|X^{2}-X^{1}\|^{k}_{\gamma}]$
which follows from independent coupling of the marginal laws $\mathcal{L}(X^{i})=\mu^{i},i=1,2.$
Remark 5.4.
An example where condition (53) is trivially satisfied is given in the case of time-independent coefficients $A^{1,i}_{t}\equiv A^{i},C^{1,i}_{t}\equiv C^{i}$ for all $t$ and for $i=1,2.$ In this case, the last two terms on the right-hand side of the stability estimate (52) vanish.
6 Randomization
[…]
Now we are going to lift $W^{\perp}$ to a geometric rough path. To do that we define for each $(s,t)\in\Delta$, the Stratonovich integral
$(\mathbb{W}^{\perp}_{s,t})^{\mathrm{Strato}}:=\int_{s}^{t}\delta W^{\perp}_{s,r}\circ\mathrm{d}W^{\perp}_{r}$
It is well known (see [31, Chapter 3]) that for all $\omega^{\prime\prime}$ outside a $\mathbb{P}^{\prime\prime}$-null set $N^{\prime\prime}$ of $\Omega^{\prime\prime}$, the Brownian rough path
$\textbf{W}^{\perp}(\omega^{\prime\prime}):=\big(W^{\perp}(\omega^{\prime\prime}),(\mathbb{W}^{\perp})^{\mathrm{Strato}}(\omega^{\prime\prime})\big)$
belongs to the space $\mathscr{C}^{0,\gamma}_{g}$ for all $\gamma\in(1/3,1/2)$ and that the following lifting map is $(\mathcal{F}^{\perp}_{t})_{t\in\mathbb{R}_{+}}$-prog. measurable:
$\begin{array}[]{llll}\textbf{W}^{\perp}:&[0,T]\times\Omega^{\prime\prime}&\rightarrow&(\mathscr{C}^{0,\gamma}_{g},\mathcal{B}(\mathscr{C}^{0,\gamma}_{g}))\\ &(t,\omega^{\prime\prime})&\mapsto&\textbf{W}^{\perp}(\omega^{\prime\prime})_{.\wedge t}.\end{array}$
Just to be clear, for $\omega^{\prime\prime}\in N^{\prime\prime}$ we set $\textbf{W}^{\perp}(\omega^{\prime\prime}):=(0,0)$, and the same convention will hold for every new process defined below whenever we randomize with respect to $\textbf{W}^{\perp}$.
Remark 6.1.
When we write $\mathbb{E},\mathbb{E}^{\prime},\mathbb{E}^{\prime\prime}$ we mean that we take expectations with respect to $\mathbb{P},\mathbb{P}^{\prime},\mathbb{P}^{\prime\prime}$.
Next, we define the following spaces of controls and vector flows.
Definition 6.2.
Let $p,d\in\mathbb{N},\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$.
$\mathcal{A}$ $:=\big\{\alpha:\Omega^{\prime}\times[0,T]\rightarrow\mathbb{R}^{\kappa}:\text{prog. measurable with respect to}\hskip 2.84544pt$
$\hskip 65.44142pt(\mathcal{F}_{t})_{t\in[0,T]},\hskip 2.84544pt\text{and}\hskip 2.84544pt\sup_{t\in[0,T]}\mathbb{E}^{\prime}[|\alpha_{t}|^{k}]<\infty,\hskip 2.84544pt\text{for every}\hskip 2.84544ptk\in[2,\infty)\big\},$
$\mathcal{A}^{1}$ $:=\big\{\alpha:\Omega^{\prime}\times\mathscr{C}^{0,\gamma}_{g}\times[0,T]\rightarrow\mathbb{R}^{\kappa}:\text{prog. measurable with respect to}\hskip 2.84544pt$
$\hskip 65.44142pt\big(\mathcal{F}_{t}\otimes\{\emptyset,\mathscr{C}^{0,\gamma}_{g}\}\big)_{t\in[0,T]},\hskip 2.84544pt\text{and}\hskip 2.84544pt{\alpha}(\boldsymbol{\eta})\in\mathcal{A}\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}\big\},$
$\mathcal{A}^{2}$ $:=\big\{\alpha:\Omega^{\prime}\times\mathscr{C}^{0,\gamma}_{g}\times[0,T]\rightarrow\mathbb{R}^{\kappa}:\text{prog. measurable with respect to}\hskip 2.84544pt$
$\hskip 65.44142pt\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]},\hskip 2.84544pt\text{and}\hskip 2.84544pt{\alpha}(\boldsymbol{\eta})\in\mathcal{A}\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}\big\},$
$\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ $:=\big\{(m,m^{\prime}):\big[m:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta}\big([0,T];\mathbb{R}^{d}\big)\hskip 2.84544pt\text{measurable with respect to}\hskip 2.84544pt\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\hskip 2.84544pt$
$\hskip 36.98866pt\text{and}\hskip 2.84544ptm^{\prime}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta^{\prime}}\big([0,T];\mathbb{R}^{d\times p}\big)\hskip 2.84544pt\text{measurable with respect to}\hskip 2.84544pt\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big),\hskip 2.84544pt$
$\hskip 36.98866pt\text{such that}\hskip 2.84544pt(m(\boldsymbol{\eta}),m^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}\big]\big\}.$
Remark 6.3.
One can also define the following space of $\boldsymbol{\eta}$-causal controls
$\mathcal{A}^{\text{Causal}}$ $:=\big\{\alpha:\Omega^{\prime}\times\mathscr{C}^{0,\gamma}_{g}\times[0,T]\rightarrow\mathbb{R}^{\kappa}:\text{prog. measurable with respect to}\hskip 2.84544pt$
$\hskip 19.91684pt\big(\mathcal{F}_{t}\otimes\mathcal{B}(\mathscr{C}^{0,\gamma}_{g})\big)_{t\in[0,T]},{\alpha}(\boldsymbol{\eta})\in\mathcal{A}\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g},\hskip 2.84544pt\text{and}\hskip 2.84544pt\alpha_{t}(\omega^{\prime},\boldsymbol{\eta})=\alpha_{t}(\omega^{\prime},\boldsymbol{\eta}_{\cdot\wedge t})\big\}.$
Note that $\mathcal{A}^{1}\subseteq\mathcal{A}^{\text{Causal}}\subseteq\mathcal{A}^{2}$, and the randomized versions of controls $\alpha\in\mathcal{A}^{\text{Causal}}$, $\overline{\alpha}$, are $(\mathcal{F}_{t}\vee\mathcal{F}^{\perp}_{t})_{t\in\mathbb{R}_{+}}$ prog. measurable processes.
At this point we introduce the following assumption.
  1. (S6-R)
    There exist $A^{1},C^{1}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$ and $(A^{1})^{\prime},(C^{1})^{\prime}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta^{\prime}}\big([0,T];L\big(\mathbb{R}^{p};L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$, such that $(A^{1}(\boldsymbol{\eta}),(A^{1})^{\prime}(\boldsymbol{\eta})),(C^{1}(\boldsymbol{\eta}),(C^{1})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, and
    $\|(A^{1},(A^{1})^{\prime});(A^{1},(A^{1})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}+\|(C^{1},(C^{1})^{\prime});(C^{1},(C^{1})^{\prime})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1}),$
    for every $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}$. Furthermore, $(C^{1}(\boldsymbol{\eta}),(C^{1})^{\prime}(\boldsymbol{\eta}))$ satisfies assumption (S4-R) for every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$.
Remark 6.4.
  1. (i)
    Following Remark 4.11, one can alternatively make the following assumption with respect to $(C^{1},(C^{1})^{\prime})$.
    1. (S7-R)
      There exist $\lambda^{(2)}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta}\big([0,T];\mathbb{R}^{1\times p}\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$ and $(\lambda^{(2)})^{\prime}:\mathscr{C}^{0,\gamma}_{g}\rightarrow C^{\beta^{\prime}}\big([0,T];L(\mathbb{R}^{p},\mathbb{R}^{1\times p})\big)$ measurable with respect to $\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)$, such that $(\lambda^{(2)}(\boldsymbol{\eta}),$ $(\lambda^{(2)})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{1\times p}\big)\hskip 2.84544pt\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, and
      $\|(\lambda^{(2)}(\boldsymbol{\eta}_{1}),(\lambda^{(2)})^{\prime}(\boldsymbol{\eta}_{1}));(\lambda^{(2)}(\boldsymbol{\eta}_{2}),(\lambda^{(2)})^{\prime}(\boldsymbol{\eta}_{2}))\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1}),$
      for every $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}$.
  2. (ii)
    From (0S7-R) we immediately get that
    $(\lambda^{(2)}(\boldsymbol{\eta})\text{Id}_{d\times d},(\lambda^{(2)})^{\prime}(\boldsymbol{\eta})\text{Id}_{d\times d})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$
    $\text{for every}\hskip 2.84544pt\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, and
    $\|(\lambda^{(2)}(\boldsymbol{\eta}_{1})\text{Id}_{d\times d},(\lambda^{(2)})^{\prime}(\boldsymbol{\eta}_{1})\text{Id}_{d\times d});(\lambda^{(2)}(\boldsymbol{\eta}_{2})\text{Id}_{d\times d},(\lambda^{(2)})^{\prime}(\boldsymbol{\eta}_{2})\text{Id}_{d\times d})\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\lesssim\rho_{\gamma}(\boldsymbol{\eta}_{2},\boldsymbol{\eta}_{1}),$
    for every $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}$.
  3. (iii)
    An example when assumption (S6-R) is satisfied is given in the case that $A^{1},C^{1}$ take values in $C^{\beta+\beta^{\prime}}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ and are constant with respect to $\mathscr{C}^{0,\gamma}_{g}$. Then, $(A^{1},0),$ $(C^{1},0)\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$ for every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$. For more examples one can apply the results of [31, Chapter 7].
Now, let $\xi\in\mathbb{R}^{d}$, $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p})$, $\alpha\in\mathcal{A}^{2}$ and $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$. Then, under S6-R, from 2.9 and [42, Proposition 3.10], we get that there exists a unique prog. measurable with respect to $\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]}$ process $X^{\boldsymbol{\eta}}_{t}(\omega^{\prime}):=X^{\boldsymbol{\eta},\alpha(\boldsymbol{\eta}),(m(\boldsymbol{\eta}),m^{\prime}(\boldsymbol{\eta})),\xi}_{t}(\omega^{\prime})$ which solves
$X^{\boldsymbol{\eta}}_{t}=\xi+\int_{0}^{t}\big(A_{s}X^{\boldsymbol{\eta}}_{s}+B_{s}\alpha(\boldsymbol{\eta})_{s}+C_{s}m(\boldsymbol{\eta})_{s}\big)\mathrm{d}s+\int_{0}^{t}\Sigma_{s}\mathrm{d}W_{s}+\int_{0}^{t}\big(A^{1}_{s}X^{\boldsymbol{\eta}}_{s}+C^{1}_{s}m(\boldsymbol{\eta})_{s}\big)\mathrm{d}\boldsymbol{\eta}_{s},$
with $(X^{\boldsymbol{\eta}},A^{1}X^{\boldsymbol{\eta}}+C^{1}m(\boldsymbol{\eta}))\in\bigcap_{k=2}^{\infty}\textbf{D}^{\gamma,\beta}_{\eta}\mathbb{L}^{k}\big([0,T];\Omega^{\prime};\mathbb{R}^{d}\big)$.
For every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, under (S6-R), there exist unique $(m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta}))\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta}\big([0,T];\mathbb{R}^{d}\big)$ and $\alpha^{*}(\boldsymbol{\eta})\in\mathcal{A}$ such that
$\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m^{*}(\boldsymbol{\eta})}_{t}\big({X}^{*,\boldsymbol{\eta}}_{t},\alpha^{*}(\boldsymbol{\eta})_{t}\big)\mathrm{d}t+\Phi^{m^{*}(\boldsymbol{\eta})}\big({X}^{*,\boldsymbol{\eta}}_{T}\big)\right]=\textrm{min}_{\alpha\in\mathcal{A}}\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m^{*}(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta},\alpha}_{t},\alpha_{t}\big)\mathrm{d}t+\Phi^{m^{*}(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta},\alpha}_{T}\big)\right]$
where
$X^{*,\boldsymbol{\eta}}_{t}(\omega^{\prime}):=X^{\boldsymbol{\eta},\alpha^{*}(\boldsymbol{\eta}),(m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta})),\xi}_{t}(\omega^{\prime}),\hskip 5.69046ptX^{\boldsymbol{\eta},\alpha}_{t}(\omega^{\prime}):=X^{\boldsymbol{\eta},\alpha,(m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta})),\xi}_{t}(\omega^{\prime}),$
and
$m^{*}(\boldsymbol{\eta})_{t}=\mathbb{E}^{\prime}\big[X^{*,\boldsymbol{\eta}}_{t}\big],\hskip 5.69046pt(m^{*})^{\prime}(\boldsymbol{\eta})_{t}=(A^{1}(\boldsymbol{\eta})_{t}+C^{1}(\boldsymbol{\eta})_{t})\mathbb{E}^{\prime}\big[X^{*,\boldsymbol{\eta}}_{t}\big].$
From Lemma 5.2, we have that $m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta})$ depend continuously on $\boldsymbol{\eta}$, thus the decoupling coefficients $\mathcal{K}^{\boldsymbol{\eta}}(\cdot,\cdot),M^{\boldsymbol{\eta}}\big((m^{*}(\boldsymbol{\eta}),(m^{*})^{\prime}(\boldsymbol{\eta}))\big)_{\cdot},P^{\boldsymbol{\eta}}_{\cdot},\Pi^{\boldsymbol{\eta},(m^{*}({\boldsymbol{\eta}}),(m^{*})^{\prime}(\boldsymbol{\eta}))}_{\cdot}$ also depend continuously on $\boldsymbol{\eta}$ (and are deterministic). Moreover, we remind that $\mathscr{C}^{0,\gamma}_{g}$ is Polish, hence Carathéodory’s Lemma can be applied, see for instance [3, 4.51 Lemma]. So, $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$, $X^{*,\boldsymbol{\eta}}$ is prog. measurable with respect to $\big(\mathcal{F}_{t}\otimes\mathcal{B}\big(\mathscr{C}^{0,\gamma}_{g}\big)\big)_{t\in[0,T]}$ and $\alpha^{*}\in\mathcal{A}^{2}$, hence we have proved the following:
Theorem 6.5.
Let $p,d\in\mathbb{N},\xi\in\mathbb{R}^{d},\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. Additionally, let (S2), (S3), (S6-R) hold true, $Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0$ and $Q+\bar{Q}-\bar{Q}S\succeq 0$. Then, there exist unique $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ and $\alpha^{*}\in\mathcal{A}^{2}$ such that, for every $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g}$, we have
$\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m^{*}(\boldsymbol{\eta})}_{t}\big({X}^{*,\boldsymbol{\eta}}_{t},\alpha^{*}(\boldsymbol{\eta})_{t}\big)\mathrm{d}t+\Phi^{m^{*}(\boldsymbol{\eta})}\big({X}^{*,\boldsymbol{\eta}}_{T}\big)\right]=\min_{\alpha\in\mathcal{A}}\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m^{*}(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta},\alpha}_{t},\alpha_{t}\big)\mathrm{d}t+\Phi^{m^{*}(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta},\alpha}_{T}\big)\right]$
while
$m^{*}(\boldsymbol{\eta})_{t}=\mathbb{E}^{\prime}\big[X^{*,\boldsymbol{\eta}}_{t}\big],\hskip 5.69046pt(m^{*})^{\prime}(\boldsymbol{\eta})_{t}=(A^{1}(\boldsymbol{\eta})_{t}+C^{1}(\boldsymbol{\eta})_{t})\mathbb{E}^{\prime}\big[X^{*,\boldsymbol{\eta}}_{t}\big].$
Moving on, for $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d}),\alpha\in\mathcal{A}^{2}$, we now define the following randomized versions as
$\hskip 99.58464pt\overline{\alpha}(\omega)$ $=\alpha(\boldsymbol{\eta})|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})},$
$\overline{m}(\omega)$ $=m(\boldsymbol{\eta})|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})},$
$\overline{m}^{\prime}(\omega)$ $=m^{\prime}(\boldsymbol{\eta})|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})},$
$\overline{X}^{\overline{\alpha}}_{t}(\omega)$ $=X^{\boldsymbol{\eta},\alpha(\boldsymbol{\eta}),(m(\boldsymbol{\eta}),m^{\prime}(\boldsymbol{\eta})),\xi}_{t}(\omega^{\prime})\big|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}.$
From the above [^12] See also [35, Theorem 3.6]. we get that $\overline{\alpha}\in\overline{\mathcal{A}},\hskip 2.84544pt\overline{m}\hskip 2.84544pt\text{is}\hskip 2.84544pt\mathcal{F}^{\perp}_{T}-\text{measurable},\hskip 2.84544pt\overline{m}^{\prime}\hskip 2.84544pt\text{is}\hskip 2.84544pt\mathcal{F}^{\perp}_{T}-\text{measurable}\hskip 2.84544pt,(\overline{m},\overline{m}^{\prime})\in\overline{\mathscr{M}}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})\hskip 2.84544pt\text{and}\hskip 2.84544pt\overline{X}\hskip 2.84544pt\text{is}\hskip 2.84544pt(\mathcal{F}_{t}\vee\mathcal{F}^{\perp}_{T})_{t\in\mathbb{R}_{+}}-\text{prog. measurable},$ where
$\overline{\mathcal{A}}$ $=\big\{\overline{\alpha}:\Omega^{\prime}\times\Omega^{\prime\prime}\times[0,T]\rightarrow\mathbb{R}^{\kappa}:\text{there exists}\hskip 2.84544pt{\alpha}\in\mathcal{A}^{2}$
$\hskip 56.9055pt\text{such that}\hskip 2.84544pt\overline{\alpha}_{t}(\omega^{\prime},\omega^{\prime\prime})={\alpha}_{t}(\omega^{\prime},\textbf{W}^{\perp}(\omega^{\prime\prime}))\big\},$
$\overline{\mathscr{M}}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ $=\big\{(\overline{m},\overline{m}^{\prime}):\text{there exists}\hskip 2.84544pt(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$
$\hskip 56.9055pt\text{such that}\hskip 2.84544pt(\overline{m}(\omega^{\prime\prime}),\overline{m}^{\prime}(\omega^{\prime\prime}))=\big(m\big(\textbf{W}^{\perp}(\omega^{\prime\prime})\big),m^{\prime}\big(\textbf{W}^{\perp}(\omega^{\prime\prime})\big)\big)\big\}.$
Moreover, by [34, Theorem 3.6], we get that for every $t\in[0,T]$ it holds
$\mathcal{L}(\overline{X}^{\overline{\alpha}}_{t})(\omega)=\mathcal{L}(\overline{X}^{\overline{\alpha}}_{t}|\mathcal{F}^{\perp}_{T})(\omega^{\prime\prime})=\mathcal{L}(X^{\boldsymbol{\eta}}_{t})\big|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}.$
Next, for $\boldsymbol{\eta}\in\mathscr{C}^{0,\gamma}_{g},\alpha\in\mathcal{A}^{2}$ and $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$, we have
$\ell^{m(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta}}_{t},\alpha(\boldsymbol{\eta})_{t}\big)\geq 0,$ $\hskip 5.69046pt\Phi^{m(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta}}_{T}\big)\geq 0$
$\ell^{\overline{m}}_{t}\big(\overline{X}^{\overline{\alpha}}_{t},\overline{\alpha}_{t}\big)\geq 0,$ $\hskip 5.69046pt\Phi^{\overline{m}}\big(\overline{X}^{\overline{\alpha}}_{T}\big)\geq 0$
so, by [28, Proposition 7.2] or [42, Proposition 5.2], we get for every $i\in\mathbb{N}$
$\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\Big(i\wedge\ell^{m(\boldsymbol{\eta})}_{t}\Big)\big({X}^{\boldsymbol{\eta}}_{t},\alpha(\boldsymbol{\eta})_{t}\big)\mathrm{d}t+\Big(i\wedge\Phi^{m(\boldsymbol{\eta})}\Big)\big({X}^{\boldsymbol{\eta}}_{T}\big)\right]\Bigg|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}$
$=$ $\mathbb{E}\!\left[\int_{0}^{T}\Big(i\wedge\ell^{\overline{m}}_{t}\Big)\big(\overline{X}^{\overline{\alpha}}_{t},\overline{\alpha}_{t}\big)\mathrm{d}t+\Big(i\wedge\Phi^{\overline{m}}\Big)\big(\overline{X}^{\overline{\alpha}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right].$
Hence, by the monotone convergence theorem we get
$\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta}}_{t},\alpha(\boldsymbol{\eta})_{t}\big)\mathrm{d}t+\Phi^{m(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta}}_{T}\big)\right]\Bigg|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}$ (63)
$=$ $\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m}}_{t}\big(\overline{X}^{\overline{\alpha}}_{t},\overline{\alpha}_{t}\big)\mathrm{d}t+\Phi^{\overline{m}}\big(\overline{X}^{\overline{\alpha}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right].$
We note that with a trivial lift $\mathcal{A}$ can be seen as $\mathcal{A}^{1}$. So, in case that ${a}\in\mathcal{A}$, equation (63) is written as
$\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta}}_{t},\alpha_{t}\big)\mathrm{d}t+\Phi^{m(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta}}_{T}\big)\right]\Bigg|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}$ (64)
$=$ $\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m}}_{t}\big(\overline{X}^{{\alpha}}_{t},{\alpha}_{t}\big)\mathrm{d}t+\Phi^{\overline{m}}\big(\overline{X}_{T}^{{\alpha}}\big)\bigg|\mathcal{F}^{\perp}_{T}\right].$
Before we state our final result, for $(m,m^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$, we define the following value functions.
$\mathcal{V}^{(m,m^{\prime}),1}(\omega)$ $:=\mathrm{essinf}_{\alpha\in\mathcal{A}^{1}}\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m}}_{t}\big(\overline{X}^{\overline{\alpha}}_{t},\overline{\alpha}_{t}\big)\mathrm{d}t+\Phi^{\overline{m}}\big(\overline{X}^{\overline{\alpha}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right],$
$\mathcal{V}^{(m,m^{\prime}),2}(\omega)$ $:=\mathrm{essinf}_{\alpha\in\mathcal{A}^{2}}\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m}}_{t}\big(\overline{X}^{\overline{\alpha}}_{t},\overline{\alpha}_{t}\big)\mathrm{d}t+\Phi^{\overline{m}}\big(\overline{X}^{\overline{\alpha}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right],$
$\overline{\mathcal{V}}^{(m,m^{\prime})}(\omega)$ $:=\left(\textrm{inf}_{\alpha\in\mathcal{A}}\mathbb{E}^{\prime}\!\left[\int_{0}^{T}\ell^{m(\boldsymbol{\eta})}_{t}\big({X}^{\boldsymbol{\eta}}_{t},\alpha_{t}\big)\mathrm{d}t+\Phi^{m(\boldsymbol{\eta})}\big({X}^{\boldsymbol{\eta}}_{T}\big)\right]\right)\Bigg|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}.$
Because $\mathcal{A}^{1}\subseteq\mathcal{A}^{2}$, and by the first part of the proof of [28, Theorem 7.4], we have
$\overline{\mathcal{V}}^{(m,m^{\prime})}(\omega)\leq\mathcal{V}^{(m,m^{\prime}),2}(\omega)\leq\mathcal{V}^{(m,m^{\prime}),1}(\omega),\hskip 5.69046pt\mathbb{P}-a.s..$ (65)
Let $(m^{*},(m^{*})^{\prime})\in\mathscr{M}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ , $\alpha^{*}\in\mathcal{A}^{2}$ be the objects described in 6.5. Then, by (63) and (64), we immediately get that
$\overline{\mathcal{V}}^{(m^{*},(m^{*})^{\prime})}(\omega)=\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m^{*}}}_{t}\big(\overline{X}^{\overline{\alpha}^{*}}_{t},\overline{\alpha^{*}}_{t}\big)\mathrm{d}t+\Phi^{\overline{m^{*}}}\big(\overline{X}^{\overline{\alpha^{*}}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right].$
Remark 6.6.
This is a major simplification in contrast with the situation in [28], where one does not know a priori that $\overline{\mathcal{V}}^{(m^{*},(m^{*})^{\prime})}$ is $\mathcal{F}^{\perp}_{T}$-measurable.
Thus, we have
$\overline{\mathcal{V}}^{(m^{*},(m^{*})^{\prime})}(\omega)=\mathrm{essinf}_{\alpha\in\mathcal{A}}\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m^{*}}}_{t}\big(\overline{X}^{{\alpha}}_{t},{\alpha}_{t}\big)\mathrm{d}t+\Phi^{\overline{m^{*}}}\big(\overline{X}^{{\alpha}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right],$
and by viewing again $\mathcal{A}$ as $\mathcal{A}^{1}$, we get by (65) that
$\overline{\mathcal{V}}^{(m^{*},(m^{*})^{\prime})}(\omega)=\mathcal{V}^{(m^{*},(m^{*})^{\prime}),2}(\omega)=\mathcal{V}^{(m^{*},(m^{*})^{\prime}),1}(\omega),\hskip 5.69046pt\mathbb{P}-a.s..$ (66)
Thus we arrive at the following result.
Theorem 6.7.
Let $p,d\in\mathbb{N},\xi\in\mathbb{R}^{d},\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. Additionally, let (S2), (S3), (S6-R) hold true, $Q_{t}+\bar{Q}_{t}-\bar{Q}_{t}S_{t}\succeq 0$ and $Q+\bar{Q}-\bar{Q}S\succeq 0$. Then, there exist $(\overline{m^{*}},(\overline{m^{*}})^{\prime})\in\overline{\mathscr{M}}^{\beta,\beta^{\prime}}_{\bullet}([0,T];\mathbb{R}^{d})$ and $\overline{\alpha^{*}}\in\overline{\mathcal{A}}$ such that
$\overline{\mathcal{V}}^{(m^{*},(m^{*})^{\prime})}(\omega)=\mathcal{V}^{(m^{*},(m^{*})^{\prime}),2}(\omega)=\mathcal{V}^{(m^{*},(m^{*})^{\prime}),1}(\omega)=\mathbb{E}\!\left[\int_{0}^{T}\ell^{\overline{m^{*}}}_{t}\big(\overline{X}^{\overline{\alpha^{*}}}_{t},\overline{\alpha^{*}}_{t}\big)\mathrm{d}t+\Phi^{\overline{m^{*}}}\big(\overline{X}^{\overline{\alpha^{*}}}_{T}\big)\bigg|\mathcal{F}^{\perp}_{T}\right]$
while
$\overline{m^{*}}_{t}=\mathbb{E}\Big[\overline{X}^{\overline{\alpha^{*}}}_{t}\big|\mathcal{F}^{\perp}_{T}\Big],\quad(\overline{m^{*}})^{\prime}_{t}=(A_{t}^{1}+C_{t}^{1})|_{\boldsymbol{\eta}=\textbf{W}^{\perp}(\omega^{\prime\prime})}\mathbb{E}\Big[\overline{X}^{\overline{\alpha^{*}}}_{t}\big|\mathcal{F}^{\perp}_{T}\Big].$
Appendix A Appendix
A.1 Linear RDEs
Proposition A.1.
Let $p,d\in\mathbb{N},\gamma\in(\frac{1}{3},\frac{1}{2})$ and $0<\beta^{\prime}\leq\beta\leq\gamma$ such that $\gamma+\beta+\beta^{\prime}>1$. Additionally, let $\xi_{1},\xi_{2}\in\mathbb{R}^{d}$ and $\boldsymbol{\eta}_{1},\boldsymbol{\eta}_{2}\in\mathscr{C}^{0,\gamma}_{g}([0,T];\mathbb{R}^{p}).$ For $T>0,$ $i=1,2$ and given $B_{i}\in C([0,T];\mathbb{R}^{d}),$ $(A_{i},A_{i}^{\prime})\in\mathscr{D}^{\beta,\beta^{\prime}}_{\eta_{i}}\big([0,T];L(\mathbb{R}^{d},\mathbb{R}^{d\times p})\big)$, there is a unique (global) solution $(X_{i},A_{i}X_{i})$ to
$dX_{i}(t)=B_{i}(t)X_{i}(t)\mathrm{d}t+(A_{i},A_{i}^{\prime})(t)X_{i}(t)d\boldsymbol{\eta}_{i},X_{i}(0)=\xi_{i}$
on $[0,T].$ Then for a constant $M>0$ such that
$\rho_{\gamma}(\boldsymbol{\eta}_{1})+\rho_{\gamma}(\boldsymbol{\eta}_{2})+\|B_{1}\|_{\infty}+\|B_{2}\|_{\infty}+\|A_{2},A_{2}^{\prime}\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}+\|A_{1},A_{1}^{\prime}\|_{\eta_{1},\eta_{2},\beta,\beta^{\prime}}\leq M$
we have